### v1

Abstract.
Let $X$ be a metric space with bounded geometry. We show that if $X$ admits a coarse embedding into an $\ell^{p}$-space ($1\leq p<\infty$), then the canonical inclusion from any geometric ideal to the corresponding ghostly ideal induces an isomorphism in $K$-theory. Our approach relies on the construction of twisted Roe algebras, avoiding the use of groupoid techniques. As consequences, we deduce the relative (maximal) coarse Baum-Connes conjectures for such spaces, as well as the Operator Norm Localization property for finite rank projections ($ONL_{\mathcal{P}_{Fin}}$).
1. Introduction
Ghosts are among the most mysterious operators in the Roe algebra. Let $X$ be a metric space with bounded geometry and fixed base point $x_{0}$. We denote $C^{*}(X)$ the Roe algebra of $X$ (2.2). An operator $T\in C^{*}(X)$ can be written as a $X$-by-$X$ matrix. Such $T$ is called a ghost if its matrix entries vanish at infinity; that is, $T_{xy}\to 0$ as $x,y\to\infty$. However, despite this local vanishing behavior, the operator $T$ is not necessarily compact. In other words, while the entries fade locally, the operator norm $\|\chi_{K_{R}}T\chi_{K_{R}}\|$ (where $K_{R}=B(x_{0},R)$) may not tend to zero as $R\to\infty$. Consequently, a ghost is an object that is elusive locally at infinity but remains capturable globally. The collection of all such operators forms a closed ideal in the Roe algebra, known as the ghostly ideal. It is also widely known as the ghost ideal in the literature. For a detailed discussion on the construction of ghost projections and the structural analysis of ghostly ideals, we refer the reader to [CW2004, WYI2012, KLVZ2021, LSZ2023, WZ2023].
Ghostly ideals play a crucial role in the study of the coarse Baum-Connes conjecture by providing counterexamples. Let $X$ be a metric space with bounded geometry. There is a coarse assembly map given by
$\mu:\lim_{d\to\infty}K_{*}(P_{d}(X))\to K_{*}(C^{*}(X)).$
The coarse Baum-Connes conjecture states that this map is an isomorphism. However, if $X$ is a sequence of expander graphs, there exists a ghost projection in the Roe algebra $C^{*}(X)$. This projection defines an element in $K_{0}(C^{*}(X))$ that lies outside the image of the assembly map $\mu$, as shown in [HLS2002, WYI2012].
On the other hand, the study of ideal structures in Roe algebras is also fundamental for $K$-theoretic computations. In [HRY1993], N. Higson, J. Roe, and G. Yu initially introduced geometric ideals associated with subspaces. They utilized this structure to establish the coarse Mayer-Vietoris principle, which allows for the reduction of $K$-theory calculations to subspaces. In [CW2004], X. Chen and Q. Wang extended the concept of geometric ideals from subspaces to invariant open sets in the coarse groupoid (see 2.5). They proved that every ideal in the uniform Roe algebra corresponds to an invariant open subset of the coarse groupoid. Furthermore, they proved that if the metric space $X$ has Yu’s Property A, then every ideal in the uniform Roe algebra is geometric. Later in [RW2014], J. Roe and R. Willett proved the converse, establishing that every ideal of the uniform Roe algebra is geometric if and only if the space has Property A.
More recently, Q. Wang and J. Zhang [WZ2023] introduced a refined concept of ghostly ideals associated with specific invariant open sets (see 2.5). Roughly speaking, for an invariant open set $U$, the corresponding ghostly ideal consists of operators whose matrix entries $T_{xy}$ vanish as $(x,y)$ tends to infinity along the direction of the complement $U^{c}$, rather than requiring them to vanish in all directions at infinity. When the subspace is bounded, the corresponding geometric ideal coincides exactly with the algebra of compact operators, and the corresponding ghostly ideal coincides with the original ghostly ideal of all ghost operators.
Although the requirement for these two ideals to be isomorphic at the algebraic level is very stringent (specifically, Property A), the conditions for their isomorphism at the $K$-theory level are much less restrictive. In [FS2014], M. Finn-Sell proved that if a space $X$ admits a coarse embedding into a Hilbert space, the inclusion map from the compact operator algebra to the ghost ideal induces an isomorphism on $K$-theory. In [WZ2023, WFZ2025], this result was generalized to the embedding from general geometric ideals to ghostly ideals.
The coarse Baum-Connes conjecture is closely related to coarse embeddings into Hilbert spaces. G. Yu’s milestone result proved that if a metric space with bounded geometry admits a coarse embedding into Hilbert space, then the coarse Baum-Connes conjecture holds for it. This condition implies the a-T-menability of the coarse groupoid $G(X)$, see [STY2002]. Indeed, previous results regarding the $K$-theory isomorphism between geometric and ghostly ideals rely heavily on the a-T-menability of the coarse groupoid.
Recently, the connection between the coarse Baum-Connes conjecture and the geometry of $\ell^{p}$-spaces ($1\leq p<\infty$) has become increasingly tight. In [WXYZ2024], it proves that if a space coarsely embeds into an $\ell^{p}$-space, then the coarse Baum-Connes conjecture holds for $X$. Furthermore, it is shown that the coarse Novikov conjecture holds if $X$ admits a fibred coarse embedding into an $\ell^{p}$-space in [GLWZ2024]. A natural question arises: does the $K$-theory isomorphism between geometric and ghostly ideals persist for spaces coarsely embeddable into $\ell^{p}$-spaces? In this paper, we prove the following result:
Theorem 1.1 (2.6 & 3.16).
Let $X$ be a metric space with bounded geometry. If $X$ admits a coarse embedding into some $\ell^{p}$-space (or $L^{p}$-space) with $p\in[1,\infty)$, then for any invariant open subset $U$, the inclusion $i:\mathcal{I}(X,U)\to\mathcal{G}(X,U)$ induces an isomorphism on $K$-theory, i.e.,
$i_{*}:K_{*}(\mathcal{I}(X,U))\to K_{*}(\mathcal{G}(X,U))$
is an isomorphism.
We emphasize that a-T-menability is linked to Hilbert space geometry. For a space $X$ that coarsely embeds into an $\ell^{p}$-space, its coarse groupoid is not necessarily a-T-menable. Consequently, our approach differs essentially from the methods used in [WZ2023, WFZ2025]. Inspired by [WXYZ2024], we employ the Dirac-dual-Dirac construction to reduce the problem to the study of ideals in twisted algebras. At the level of twisted algebras, the structure readily implies the $K$-theory isomorphism between these two ideals.
As applications of our main result, we derive the following corollaries. The first corollary concerns the relative coarse Baum-Connes conjecture, introduced in [GWZ2025]. Its validity characterizes local obstructions to positive scalar curvature at infinity for manifolds.
Corollary 1.2 (4.1).
Let $X$ be a metric space with bounded geometry that admits a coarse embedding into an $\ell^{p}$-space ($p\geq 1$). For any subspace $Y$, the relative coarse Baum-Connes conjecture holds for $(X,Y)$.
The second concerns the operator norm localization for equi-approximable finite-rank projections, a concept introduced in [BFV2024] that is of interest in geometric group theory. In this paper, we establish a connection between this condition and the geometry of $\ell^{p}$-spaces via $K$-theory computations.
Corollary 1.3 (4.3).
Let $X$ be a metric space with bounded geometry. If $X$ admits a coarse embedding into $\ell^{p}$, then $X$ has operator norm localization for equi-approximable finite-rank projections.
Finally, we extend our results from the reduced setting to the maximal setting. By combining our main theorem with the techniques in [HIT2020, WXYZ2024], we deduce the following result regarding the maximal coarse Baum-Connes conjecture:
Corollary 1.4 (4.8).
Let $X$ be a metric space with bounded geometry that admits a coarse embedding into an $\ell^{p}$-space ($p\geq 1$).
  • (1)
    The canonical quotient map $\pi:C^{*}_{\max}(X)\to C^{*}(X)$ induces an isomorphism on $K$-theory, i.e.,
    $\pi_{*}:K_{*}(C^{*}_{\max}(X))\xrightarrow{\cong}K_{*}(C^{*}(X)).$
  • (2)
    For any subspace $Y$, the canonical quotient map $\pi:C^{*}_{\max,Y,\infty}(X)\to C^{*}_{Y,\infty}(X)$ induces an isomorphism on the level of $K$-theory, i.e.,
    $\pi_{*}:K_{*}(C^{*}_{\max,Y,\infty}(X))\xrightarrow{\cong}K_{*}(C^{*}_{Y,\infty}(X)).$
  • (3)
    For any subspace $Y$, the maximal relative coarse Baum-Connes conjecture holds for $(X,Y)$.
Our results naturally motivate further inquiries regarding more general coefficients and embedding conditions. We propose the following open questions:
Question 1.5.
In [DG2024], a twisted version of the coarse Baum-Connes conjecture is introduced. It is natural to ask whether the twisted coarse Baum-Connes conjecture holds for $X$ with any twisted coefficient algebra if $X$ admits a coarse embedding into $\ell^{p}$-spaces?
Furthermore, parallel with [CWY2013], does the maximal coarse Baum-Connes conjecture hold for $X$ if $X$ admits a fibred coarse embedding into $\ell^{p}$-space, with $p\in[1,\infty)$?
The paper is organized as follows. In Section 2, we recall some preliminaries on coarse geometry, and ideal structures of Roe algebras. Section 3 is devoted to the proof of our main theorem in four steps. Finally, in Section 4, we discuss applications to the relative and maximal coarse Baum-Connes conjectures, as well as the operator norm localization property.
2. Geometric and ghostly ideals in Roe algebras
In this section, we shall recall some basic facts on coarse geometry and the ideal structure of Roe algebras.
2.1. Coarse groupoid
Throughout this paper, we shall always assume $(X,d)$ to be a metric space with bounded geometry, i.e., for any $R>0$,
$\sup_{x\in X}\#B(x,R)<\infty,$
where $B(x,R)=\{x^{\prime}\in X\mid d(x,x^{\prime})\leq R\}$. Denote $\beta X$ the Stone-Čech compactification of $X$, which is the set of all ultrafilters on $X$ endowed with a topology making it a compact Hausdorff space such that $C(\beta X)\cong\ell^{\infty}(X)$. The reader is referred to [Goldbring2022] for details of ultrafilters and Stone-Čech compactification. For each $R>0$, we define the $R$-diagonal of $X\times X$ by
$\Delta_{R}=\{(x,y)\in X\times X\mid d(x,y)\leq R\}\subseteq X\times X.$
When $R=0$, the set $\Delta_{0}$ is exactly the diagonal in $X\times X$. A subset of an $R$-diagonal is called an entourage (or a controlled set). The set of all entourages, denoted by $\mathcal{E}_{d}$, is called the coarse structure of $X$ associated with the metric. The coarse groupoid of $X$, denoted by $G(X)$, is defined to be
$G(X)=\bigcup_{R\geq 0}\overline{\Delta_{R}}\subseteq\beta(X\times X).$
The space $\beta X$ is the unit space of $G(X)$. The projection onto the first (second, resp.) coordinate $r:X\times X\to X$, ($s:X\times X\to X$, resp.) extends continuously to the range map $r:G(X)\to\beta X$ (the source map $s:G(X)\to\beta X$, resp.). The reader is referred to [STY2002] for details of the coarse groupoid. A subset $U\subset\beta X$ is called invariant if $r(\gamma)\in U$ is equivalent to $s(\gamma)\in U$ for any $\gamma\in G(X)$. For a subset $Y\subseteq X$, we define
$U_{Y}=\bigcup_{R\geq 0}\overline{B(Y,R)}\subseteq\beta X.$
It is direct to see that $U_{Y}$ is an invariant open set. We shall denote $F_{Y}=U_{Y}^{c}$, which is an invariant closed set.
A map $t:D\to R$ is called a partial translation in $X$ if
  • (1)
    $D,R$ are subsets of $X$ and $t$ is a bijection between $D$ and $R$;
  • (2)
    there exists $M>0$ such that $\sup_{x\in D}d(x,t(x))\leq M$.
It is proved in [SpaWil2017, Appendix C] that for any $\gamma\in G(X)$ with $s(\gamma)=\alpha\in\beta X$, there exists a partial translation $t:D\to R$ in $X$ such that $\gamma=\lim_{x\to\alpha}(t(x),x)\in G(X)$.
For any map $f:X\to Y$, we shall still denote $f$ the extended continuous map $\beta X\to\beta Y$ with a little abuse of notation. For any subset $A\subseteq\beta X$, we denote $G(X)|^{A}_{A}=r^{-1}(A)\cap s^{-1}(A)$, which is subgroupoid. Let $Y\subseteq X$ be a subset. Then the canonical inclusion map $i:Y\to X$ extends to $i:\beta Y\to\beta X$. Thus, one can identify $\beta Y$ to the closure of $Y$ in $\beta X$. As a result, we have that $G(Y)=G(X)|^{\beta Y}_{\beta Y}$. For any invariant subset $U\subseteq\beta X$, we shall denote $U^{\beta Y}=\beta Y\cap U$ and it is direct to see that $U^{\beta Y}$ is $G(Y)$-invariant. A coarse ideal pair $(X,U)$ includes a metric space $X$ and an invariant subset $U$. Two pairs $(X,U)$ and $(Y,V)$ are coarsely equivalent if there exists a coarse equivalence $f:X\to Y$ such that $V$ is the smallest invariant open set containing $f(U)$. In this case, it is equivalent to say that $V$ is the minimal invariant open subset such that $U=f^{-1}(V)$.
Examples 2.1.
Let $Y\subseteq X$ be a subset. We shall denote $Y_{R}=B(Y,R)$ the $R$-neighbourhood of $Y$, where $B(Y,R)=\bigcup_{y\in Y}B(y,R)$. Then the invariant open subset generated by $Y$, denoted by $U_{Y}$, is defined to be $U_{Y}=\bigcup_{R\geq 0}\overline{Y_{R}}\subseteq\beta X$. It is direct to check that $U_{Y}$ is indeed an invariant open subset.
Let $f:X\to Z$ be a coarse equivalent. Set $W$ to be the $G(Z)$-invariant open subset of $\beta Z$ generated by $f(Y)$. Then it is direct to check that $(X,U_{Y})$ is coarsely equivalent to $(Z,W)$.
2.2. Roe algebras and their ideals
Fix $\mathcal{H}$ to be an infinite-dimensional, separable Hilbert space. Denote $\mathcal{H}_{X}=\ell^{2}(X)\otimes\mathcal{H}$. Then there is a canonical representation of $C_{0}(X)$ on $\mathcal{H}_{X}$ by
$f\cdot\delta_{x}\otimes v=f(x)\delta_{x}\otimes v$
for $\delta_{x}\in\ell^{2}(X)$ and $v\in\mathcal{H}$. An operator $T\in\mathcal{B}(\mathcal{H}_{X})$ can be viewed as an $X$-by-$X$ matrix with value in $\mathcal{B}(\mathcal{H})$, where its $(x,y)$-th entry is given by:
$\mathcal{B}(\mathcal{H})\ni T_{xy}=\delta_{x}T\delta_{y}:\mathbb{C}\delta_{y}\otimes\mathcal{H}\to\mathbb{C}\delta_{x}\otimes\mathcal{H},$
where $\delta_{x},\delta_{y}\in C_{0}(X)$ are Dirac functions on $x$ and $y$, respectively.
Definition 2.2.
The algebra Roe algebra of $X$, denoted by $\mathbb{C}[X,\mathcal{H}_{X}]$ (or $\mathbb{C}[X]$ for simplicity), is a subalgebra of $\mathcal{B}(\mathcal{H}_{X})$ consisting of all operators $T$ satisfying:
  • $T$ has finite propagation (or controlled propagation), i.e., the support of $T$
    $\textup{supp}(T)=\{(x,y)\in X\times X\mid T_{xy}\neq 0\},$
    is an entourage.
  • $T_{xy}\in\mathcal{K}(\mathcal{H})$ for any $(x,y)\in X\times X$.
The completion of $\mathbb{C}[X]$ in $\mathcal{B}(\mathcal{H}_{X})$ is called the Roe algebra of $X$, denoted by $C^{*}(X,\mathcal{H}_{X})$ (or $C^{*}(X)$ for simplicity).
We should remark that the definition of the Roe algebra does not depend on the choice of $\mathcal{H}$ up to a non-canonical isomorphism. The following lemma is proved in [STY2002, Lemma 4.4], which reveals that Roe algebras can be fundamentally understood as groupoid $C^{*}$-algebras. If we replace $\mathcal{H}$ with $\mathbb{C}$, the resulting Roe algebra is called the uniform Roe algebra, denoted by $C^{*}_{u}(X)$.
Lemma 2.3.
Let X be a metric space with bounded geometry. There is a natural isomorphism
$C^{*}(X)\cong\ell^{\infty}(X,\mathcal{K}(\mathcal{H}))\rtimes_{r}G(X).\qed$
Let $\mathcal{I}$ be an ideal of $C^{*}(X)$. For any $T\in\mathcal{I}$, we define the $\varepsilon$-support of $T$ to be
$\textup{supp}_{\varepsilon}(T)=\{(x,y)\in X\times X\mid\|T_{xy}\|\geq\varepsilon\}.$
Since $T$ is in Roe algebra, it is direct to see that $\textup{supp}_{\varepsilon}(T)$ is an entourage. We define
$U(\mathcal{I})=\bigcup_{T\in\mathcal{I},\varepsilon>0}\overline{r(\textup{supp}_{\varepsilon}(T))}\subseteq\beta X.$
Since $\mathcal{I}$ is an ideal, one can prove that $U(\mathcal{I})$ is an invariant open subset, see [CW2004] for a detailed discussion.
Remark 2.4.
For ideals in Roe algebras, we can analyze them through two components:
  • Each ideal corresponds to a $G(X)$-invariant open set (determined by calculating operator supports as above);
  • After fixing the support, the coefficient part of each operator (i.e., the $\ell^{\infty}(X,\mathcal{K})$ component) further determines the rank distribution of operator entries within the ideal, as discussed in [CW2006, WFZ2025].
In this paper, we will not consider the $\ell^{\infty}(X,\mathcal{K})$ aspect. That is, we assume the ideal $\mathcal{I}$ has maximal rank distribution over its corresponding invariant open set, i.e., $\mathcal{I}$ contains all rank distributions supported on $U(\mathcal{I})$. We will not review the precise definition of rank distribution here; interested readers may refer to [CW2006, WFZ2025] for details.
Definition 2.5.
Let $U\subseteq\beta X$ be an invariant open set. The geometric ideal of $C^{*}(X)$ with respect to $U$, denoted by $\mathcal{I}(X,U)$, is defined to be the closed ideal generated by
$\mathbb{I}[X,U]=\left\{T\in\mathbb{C}[X]\ \big|\ \overline{r(\textup{supp}_{\varepsilon}(T))}\subseteq U\right\}.$
The ghostly ideal of $C^{*}_{u}(X)$ with respect to $U$, denoted by $\mathcal{I}_{\mathcal{G}}(X,U)$, is defined to be
$\mathcal{G}(X,U)=\left\{T\in C^{*}(X)\ \big|\ \overline{r(\textup{supp}_{\varepsilon}(T))}\subseteq U\right\}.$
One should notice that $\mathbb{I}[X,U]$ is dense in $\mathcal{I}(X,U)$ and is also an ideal of $\mathbb{C}[X]$. Note that the definitions of geometric ideals and ghostly ideals are very similar, but they have essential differences. The geometric ideal is constructed in the algebraic Roe algebra by taking the part where the $\varepsilon$-support set lies in $U$ and then generating an ideal, while the ghostly ideal is obtained in the completed Roe algebra by taking the part where the $\varepsilon$-support set lies in $U$. The essential difference lies in the order of taking the completion and imposing the support condition. Therefore, it is straightforward to see that the geometric ideal is a subalgebra of the ghostly ideal. Moreover, the ghostly ideal is larger than the corresponding geometric ideal, and the two ideals coincide only when the space has Yu’s Property A (see [WZ2023]).
Our main result of this paper is as follows.
Theorem 2.6.
Let $X$ be a metric space with bounded geometry. If $X$ admits a coarse embedding into some $\ell^{p}$-space with $p\in[1,\infty)$, then for any invariant open subset $U$, the inclusion $i:\mathcal{I}(X,U)\to\mathcal{G}(X,U)$ induces an isomorphism on $K$-theory, i.e.,
$i_{*}:K_{*}(\mathcal{I}(X,U))\to K_{*}(\mathcal{G}(X,U))$
is an isomorphism.
As a result of P. Nowak [Nowak2006], for $p\in[1,2)$, coarse embeddability into $\ell^{p}$ is equivalent to coarse embeddability into Hilbert space. In this case, the result is proved in [WFZ2025]. Therefore, we may assume $p>2$ and that there exists a coarse embedding $f:X\to\ell^{p}$. Without loss of generality, we shall assume $p\geq 2$ throughout this paper.
It is worth mentioning that for $q\in[2,\infty), p\ell^{p}$ admits a coarse embedding into $\ell^{q}$ if and only if $p\leq q$, see [MN2008, Theorem 1.10]. Thus, coarse embedding into $\ell^{p}$-space is stronger than coarse embedding into Hilbert space when $p>2$. In other words, for any metric space $X$ with bounded geometry, there exists $p_{X}\in[2,\infty]$ such that $X$ admits a coarse embedding into $\ell^{p}$ for any $p\geq p_{X}$. There is still no known example for a space with bounded geometry that coarsely embeds into $\ell^{p}$ for $p>2$ but can not coarsely embed into Hilbert space.
2.3. Coarse invariant of ghostly/geometric ideals
In this subsection, we shall show that the ghostly ideal is coarse invariant.
Proposition 2.7.
If $f:X\to Y$ is a coarse equivalent between the pair $(X,U)$ and $(Y,V)$, then $f$ induces an isometry $W_{f}:\mathcal{H}_{X}\to H_{Y}$ such that
$Ad_{W_{f}}:\mathcal{I}(X,U)\to\mathcal{I}(Y,V)\quad\text{and}\quad Ad_{W_{f}}:\mathcal{G}(X,U)\to\mathcal{G}(Y,V).$
In particular, $W_{f}$ can be chosen to be a unitary.
Proof.
By [HIT2020, Proposition 4.3.4 & Proposition 4.3.5], one can find a covering isometry $W_{f}:\mathcal{H}_{X}\to\mathcal{H}_{Y}$ for $f$. For discrete metric spaces with bounded geometry, the construction of $W_{f}$ will be easier. We shall briefly recall the process here.
Write $Z=f(X)$. We can take $A_{z}=f^{-1}(z)$. Choose a unitary $W_{z}:\ell^{2}(A_{z})\otimes\mathcal{H}\to\delta_{z}\otimes\mathcal{H}$. Then
$W_{f}=\oplus_{z\in Z}W_{z}:\bigoplus_{z\in Z}\ell^{2}(A_{z})\otimes\mathcal{H}=\mathcal{H}_{X}\to\ell^{2}(Z)\otimes\mathcal{H}\subseteq\mathcal{H}_{Y}$
is a covering isometry. To find a unitary, we can take a uniformly bounded Borel cover $\{B_{z}\}_{z\in Z}$ of $Y$ such that $z\in B_{z}$ for all $z\in Z$. Then $\{f^{-1}(B_{z})=f^{-1}(z)\}_{z\in Z}$ is also a uniformly bounded Borel cover of $X$. Write $A_{z}=f^{-1}(z)$ for simplicity. For any $z\in Z$, take an arbitrary unitary $W_{z}:\ell^{2}(A_{z})\otimes\mathcal{H}\to\ell^{2}(B_{z})\otimes\mathcal{H}$. Then
$W_{f}=\oplus_{z\in Z}W_{z}:\bigoplus_{z\in Z}\ell^{2}(A_{z})\otimes\mathcal{H}=\mathcal{H}_{X}\to\bigoplus_{z\in Z}\ell^{2}(B_{z})\otimes\mathcal{H}=\mathcal{H}_{Y}$
is exactly the covering unitary we need.
For the geometric ideal, it suffices to prove $Ad_{W_{f}}(T)\in\mathbb{I}[Y,V]$ for any $T\in\mathbb{I}[X,U]$. By the bounded geometry condition of $X$, there exists $N>0$ such that $(f^{-1}(y))\leq N$. Fix $\varepsilon>0$. Let $A\in M_{N}(\mathcal{K})$ be an NN-by-NN matrix all of whose entries have norm smaller than $\varepsilon$. Then the norm of $A$ is smaller than $\varepsilon\cdot N$. For any $y\in Y$, we shall write $z_{y}\in Z$ such that $y\in B_{z_{y}}$. From the construction of $W_{f}$, we have that
$\|Ad_{W_{f}}(T)_{yy^{\prime}}\|\leq N\cdot\sup\{\|T_{xx^{\prime}}\|\mid x\in A_{z_{y}},x^{\prime}\in A_{z_{y^{\prime}}}\}.$
Thus, if $\|Ad_{W_{f}}(T)_{yy^{\prime}}\|\geq\varepsilon$, there exists at least one pair $(x,x^{\prime})\in A_{z_{y}}\times A_{z_{y^{\prime}}}$ such that $\|T_{xx^{\prime}}\|\geq\varepsilon/N$. Denote $R=\sup_{z\in Z}\textup{diam}(A_{z})$. We then conclude that
$f^{-1}({r(\textup{supp}_{\varepsilon}(Ad_{W_{f}}(T)))})\subseteq B(r(\textup{supp}_{\varepsilon/N}(T)),R).$
By the definition of an invariant subset, we have $\overline{B(r(\textup{supp}_{\varepsilon/N}(T)),R)}\subseteq U$ since $\overline{r(\textup{supp}_{\varepsilon/N}(T))}\subseteq U$. As a result, $\overline{r(\textup{supp}_{\varepsilon}(Ad_{W_{f}}(T)))}\subseteq f(U)\subseteq V$. We conclude that $Ad_{W_{f}}(T)\in\mathbb{I}[Y,V]$. For the ghostly ideal, the proof is parallel, omitted. ∎
Corollary 2.8.
If $f:X\to Y$ is a coarse equivalent between the pair $(X,U)$ and $(Y,V)$, then
$(Ad_{W_{f}})_{*}:K_{*}(\mathcal{I}(X,U))\to K_{*}(\mathcal{I}(Y,V))\quad\text{and}\quad(Ad_{W_{f}})_{*}:K_{*}(\mathcal{G}(X,U))\to K_{*}(\mathcal{G}(Y,V))$
does not depend on the choice of $W_{f}$ that covers $f$. As a result, $(Ad_{W_{f}})_{*}$ is always an isomorphism.
Proof.
For any two covering isometry $W_{1},W_{2}$ for $f$, it suffices to prove that the *-homomorphism from $\mathcal{I}(X,U)$ to $M_{2}(\mathcal{I}(Y,V))$ defined by
$\alpha_{1}:T\mapsto\begin{pmatrix}W_{1}TW_{1}^{*}&0\\ 0&0\end{pmatrix}\quad\text{and}\quad\alpha_{2}:T\mapsto\begin{pmatrix}0&0\\ 0&W_{2}TW_{2}^{*}\end{pmatrix}$
agree on the level of $K$-theory. Take the unitary
$U=\begin{pmatrix}1-V_{1}V_{1}^{*}&V_{1}V_{2}^{*}\\ V_{2}V_{1}^{*}&1-V_{2}V_{2}^{*}\end{pmatrix}$.
Conjugation by $U$ induces the identity map on $K$-theory and one can easily chech that $U\alpha_{1}(T)U^{*}=\alpha_{2}(T)$. This completes the proof.
Moreover, $W_{f}$ can be chosen to be a unitary, in which case $Ad_{W_{f}}$ is a *-isomorphism. This proves that $(Ad_{W_{f}})_{*}$ is always an isomorphism. ∎
2.4. Reduction to sparse spaces
In this subsection, we shall reduce 2.6 to the case of sparse spaces.
By [CWY2013, Section 4], one can find an $\omega$-excisive cover $\{Y,Z\}$ of $X$ such that $Y$, $Z$, and $Y\cap Z$ are all sparse, i.e., they are all coarse disjoint unions of finite spaces. From the definition of Stone-Čech compactification, we conclude that $\beta X=\beta Y\cup\beta Z$. Then the canonical inclusion $\iota:Y\to X$ induces a canonical inclusion $\iota:G(Y)\to G(X)$. Actually, one can easily check the identification $G(Y)=G(X)|^{\beta Y}_{\beta Y}$.
If $U\subseteq\beta X$ is an invariant open subset, then it is from the definition that $U^{\beta Y}=U\cap\beta Y$ and $U^{\beta Z}=U\cap\beta Z$ are $G(Y)$-invariant and $G(Z)$-invariant, respectively. In this case, we shall denote $\mathcal{I}(Y,U^{\beta Y})$ and ${\mathcal{G}}(Y,U^{\beta Y})$ the corresponding geometric and ghostly ideal of $C^{*}(Y)$ with respect to $U^{\beta Y}$. We have the following observation.
Lemma 2.9.
With notations as above, we have that
$\mathcal{I}(Y,U^{\beta Y})=\mathcal{I}(X,U)\cap C^{*}(Y)\quad\text{and}\quad\mathcal{G}(Y,U^{\beta Y})=\mathcal{G}(X,U)\cap C^{*}(Y).$
Proof.
For the geometric ideal, it suffices to show
$\mathbb{I}[Y,U^{\beta Y}]=\mathbb{I}[X,U]\cap\mathbb{C}[Y]$
since $\mathbb{I}[Y,U^{\beta Y}]$, $\mathbb{I}[X,U]$, $\mathbb{C}[Y]$ are respectively dense in $\mathcal{I}(Y,U^{\beta Y})$, $\mathcal{I}(X,U)$, $C^{*}(Y)$. For any $T\in\mathbb{I}[Y,U^{\beta Y}]$, it is direct to see that $r(\textup{supp}_{\varepsilon}(T))\subseteq Y\cap U$ from definition. Thus, $T\in\mathbb{I}[X,U]$. On the other hand, if $T\in\mathbb{C}[Y]$ satisfies that $\overline{r(\textup{supp}_{\varepsilon}(T))}\subseteq\beta Y\cap U$, then $r(\textup{supp}_{\varepsilon}(T))\subseteq Y\cap U$. This shows that $T\in\mathbb{I}[X,U]$. For the ghostly ideal, the proof is parallel, omitted. ∎
Write $V_{Y}$ the $G(X)$-invariant open subset in $\beta X$. Notice that $\mathcal{I}(X,V_{Y})$ is isomorphic to the ideal of $C^{*}(X)$ generated by $Y$, denoted by $C^{*}(X,Y)$ as in [HRY1993]. Thus, one has that
$\mathcal{I}(X,V_{Y})=\lim_{r\to\infty}C^{*}(B(Y_{R})).$
As a result, there exist canonical inclusions
$\mathcal{I}(Y,U^{\beta Y})\hookrightarrow C^{*}(Y)\hookrightarrow\mathcal{I}(X,V_{Y}).$
Lemma 2.10.
With notation as above, we have that the canonical inclusion map $\iota:C^{*}(Y)\cap\mathcal{I}\to\mathcal{I}(X,V_{Y})\cap\mathcal{I}$ induces an isomorphism on $K$-theory
$\iota_{*}:K_{*}(C^{*}(Y)\cap\mathcal{I})\xrightarrow{\cong}K_{*}(\mathcal{I}(X,V_{Y})\cap\mathcal{I}),$
where $\mathcal{I}=\mathcal{I}(X,U)$ or $\mathcal{I}=\mathcal{G}(X,U)$.
Proof.
As we discussed above, the inclusion map can be induced by the inductive limit of the inclusions as $R$ tends to infinity
$\iota_{RR^{\prime}}:C^{*}(Y_{R})\cap\mathcal{I}\to C^{*}(Y_{R^{\prime}})\cap\mathcal{I}.$
By 2.8, it induces an isomorphism on $K$-theory, i.e.,
$(\iota_{RR^{\prime}})_{*}:K_{*}(C^{*}(Y_{R})\cap\mathcal{I})\xrightarrow{\cong}K_{*}(C^{*}(Y_{R^{\prime}})\cap\mathcal{I})$
is an isomorphism for any $R,R^{\prime}\geq 0$. As a result, the inclusion $\iota$ induces an isomorphism on $K$-theory, i.e.,
$\iota_{*}:K_{*}(C^{*}(Y)\cap\mathcal{I})\xrightarrow{\cong}K_{*}(\mathcal{I}(X,V_{Y})\cap\mathcal{I}).$
This completes the proof. ∎
Proposition 2.11.
If 2.6 holds for all sparse subspaces $Y$ of $X$, then it holds for $X$ itself.
Proof.
Take $\{Y,Z\}$ to be an $\omega$-excisive cover of $X$ such that $Y,Z$ and $Y\cap Z$ are all sparse. Since $Y\cup Z=X$, we have $V_{Y}\cup V_{Z}=\beta X$ by the definition of the Stone-Čech compactification, where $V_{Y}$ and $V_{Z}$ are $G(X)$-invariant open subset generated by $Y$ and $Z$, respectively. It is proved in [HRY1993] that
$\mathcal{I}(X,V_{Y})+\mathcal{I}(X,V_{Z})=C^{*}(X)\quad\text{and}\quad\mathcal{I}(X,V_{Y})\cap\mathcal{I}(X,V_{Z})=\mathcal{I}(X,V_{Y\cap Z}).$
The second equation follows from the fact that $\{Y,Z\}$ is $\omega$-excisive. It is direct to see that $V_{Y}\cap V_{Z}=V_{Y\cap Z}$.
For an invariant open subset $U\subseteq\beta X$, let $\mathcal{I}$ denote either the geometric ideal $\mathcal{I}(X,U)$ or the ghostly ideal $\mathcal{G}(X,U)$. By taking intersections with $\mathcal{I}$ in the pushout diagram of the geometric and ghostly ideals mentioned above, we obtain the following pushout diagram:
$\mathcal{I}(X,U_{Y\cap Z})\cap\mathcal{I}$$\mathcal{I}(X,U_{Z})\cap\mathcal{I}$$\mathcal{I}(X,U_{Y})\cap\mathcal{I}$$\mathcal{I}$}
By the Five Lemma, it suffices to prove the canonical inclusion map
$i:\mathcal{I}(X,U_{W})\cap\mathcal{I}(X,U)\to\mathcal{I}(X,U_{W})\cap\mathcal{G}(X,U)$
induces an isomorphism on $K$-theory for $W=Y,Z$ and $Y\cap Z$ which are all sparse. By 2.10, we conclude that $i_{*}$ is identified with
$i_{*}:K_{*}(C^{*}(W)\cap\mathcal{I}(X,U))\to K_{*}(C^{*}(W)\cap\mathcal{G}(X,U)).$
By 2.9, the map $i_{*}$ is further identified with
$i_{*}:K_{*}(\mathcal{I}(W,U^{\beta W}))\to K_{*}(\mathcal{G}(W,U^{\beta W})).$
Therefore, it suffices to prove 2.6 for the case when $X$ is a sparse space, as the general case will then follow. This completes the proof of the proposition. ∎
Remark 2.12.
For ghostly ideals, we actually have the parallel pushout diagram, i.e.,
$\mathcal{G}(X,V_{Y})+\mathcal{G}(X,V_{Z})=C^{*}(X)\quad\text{and}\quad\mathcal{G}(X,V_{Y})\cap\mathcal{G}(X,V_{Z})=\mathcal{G}(X,V_{Y\cap Z}).$
Since $\mathcal{I}(X,V_{Y})\subseteq\mathcal{G}(X,V_{Y})$, the first equation follows from that $\mathcal{I}(X,V_{Y})+\mathcal{I}(X,V_{Z})=C^{*}(X)$. For the second one, take an element $T\in C^{*}(X)$ and an invariant open subset $V\subseteq\beta X$. Define a function $f_{T}:X\times X\to\mathbb{R}$ by
$f_{T}(x,y)=\|T_{xy}\|.$
By definition, $f_{T}$ admits an extension to a function in $C_{0}(G(X))$. The reasoning is that $T$ can be approximated by finite propagation operators. Since any such operator is supported in a set of the form $\overline{\Delta_{R}}$, its associated function lies in $C_{c}(G(X))$. It is direct from the definition that $T\in\mathcal{G}(X,V)$ if and only if $r(\textup{supp}(f_{T}))\subseteq V$. As a result, $T\in\mathcal{G}(X,V_{Y})\cap\mathcal{G}(X,V_{Z})$ if and only if
$r(\textup{supp}(f_{T}))\subseteq V_{Y}\cap V_{Z}=V_{Y\cap Z}.$
This proves the pushout diagram of ghostly ideals.
3. The $\ell^{p}$-Bott-Dirac construction and twisted ghostly ideals
In this section, we prove 2.6. The argument proceeds in the following three steps.
3.1. Step 1. On the Bott-Dirac operator for $\ell^{p}$-spaces
In the following subsections, we shall prove 2.6 in several steps. First of all, we shall briefly recall the definition of the Bott-Dirac operator in finite-dimensional $\ell^{p}$-spaces. Please refer to [WXYZ2024] for a detailed introduction.
Let $E=\mathbb{C}^{n}$ be an $n$-dimensional vector space with a canonical basis $\{e_{1},\cdots,e_{n}\}$. For any $p\geq 1$, we can equip it with the $\ell^{p}$-norm associated with the given basis $\|\cdot\|_{p}$. Denote by $S_{p}(E)$ the sphere of $E$ under the norm $\|\cdot\|_{p}$. Consider the Mazur map
$\psi:S_{p}(E)\to S_{2}(E),\quad\sum_{i=1}^{n}a_{i}e_{i}\mapsto\sum_{i=1}^{n}a_{i}|a_{i}|^{p/2-1}e_{i}.$
We can extend the Mazur map homogeneously to
$\Psi:(E,\|\cdot\|_{p})\to(E,\|\cdot\|_{2}),\quad v\mapsto\left\{\begin{aligned} &\|v\|_{p}\psi\left(\frac{v}{\|v\|_{2}}\right),&&\text{if }v\neq 0\\ &0,&&\text{if }v=0\end{aligned}\right.$.
When $p\geq 2$, the extended Mazur map is Lipschitz with constant $L$.
Denote by $\textup{Cliff}_{\mathbb{C}}(E)$ the Clifford algebra associated with $(E,\|\cdot\|_{2})$. It is easy to verify that if $E$ is an $n$-dimensional space, then $\textup{Cliff}_{\mathbb{C}}(E)$ is a vector space of dimension $2^{n}$. Given an orthonormal basis $\{e_{1},\dots,e_{n}\}$ of $E$, the set
$\{e_{i_{1}}e_{i_{2}}\cdots e_{i_{k}}\mid 1\leq i_{1}<i_{2}<\cdots<i_{k}\leq n,\ 0\leq k\leq n\},$
forms a basis for $\textup{Cliff}_{\mathbb{C}}(E)$. This basis then defines a Hilbert space structure on $\textup{Cliff}_{\mathbb{C}}(E)$, which we denote by $\mathfrak{h}_{E}$. Moreover, $\textup{Cliff}_{\mathbb{C}}(E)$ acts canonically on $\mathfrak{h}_{E}$ by left multiplication.
For any $v\in(E,\|,\|_{p})$, we define the Bott generator to be the unbounded function
$C_{v}:(E,\|\cdot\|_{p})\to\textup{Cliff}_{\mathbb{C}}(E),\quad w\mapsto\Psi(w-v)$
Define the Dirac operator $D$ to be
$D=\sum_{i=1}^{n}\widehat{e_{i}}\frac{\partial}{\partial x_{i}},$
where $\widehat{v}:\textup{Cliff}_{\mathbb{C}}(E)\to\textup{Cliff}_{\mathbb{C}}(E)$ is defined to be $\widehat{v}(w)=(-1)^{deg(w)}wv$ for any $v\in E$ and $w\in\textup{Cliff}_{\mathbb{C}}(E)$. Here, $D$ is viewed as an unbounded operator on $L^{2}(E,\mathfrak{h}_{E})$
For any $s\geq 1$, we define the Bott-Dirac operator to be
$B_{s,v}=s^{-1}D+C_{v}$
We summarize the facts we will need about the Bott-Drac operator in the next result. A detailed proof for the proposition can be founded in [WXYZ2024, Lemma 2.7]:
Proposition 3.1 ([WXYZ2024]).
For any $s>1$ and $v\in E$, the operator $B_{s,v}$ is essentially self-adjoint on $\mathfrak{h}_{E}$ and admits a compact resolvent. Furthermore, it is a Fredholm operator with index $1$.
The following bounded version of the Bott-Dirac operator will be frequently used in what follows.
Definition 3.2.
For any $v\in E$ and $s>1$, we define the bounded Bott-Dirac operator to be
$F_{s,v}:=B_{s,v}(1+B^{2}_{s,v})^{-\frac{1}{2}}.$
Notice that $F_{s,v}$ is a bounded operator on $L^{2}(E,\mathfrak{h}_{E})$. There is a natural action of $C_{0}(E)$ on $L^{2}(E,\mathfrak{h}_{E})$, given by pointwise multiplication. Consequently, for any $p\in[1,\infty)$, we can discuss the propagation of Fs,vF{s,v} in the $E$-direction with respect to the norm $|\cdot|_{p}$. We shall denote this number by $\textup{Prop}_{E,p}(F_{s,v})$.
For any $v\in E$ and $R>0$, we denote $\chi_{v,R}:=\chi_{B(v,R)}$ for simplicity. The following theorem summarizes some properties of $F_{s,v}$. For the detailed proof, we refer the reader to [WXYZ2024, Proposition 2.17].
Proposition 3.3.
For each $\varepsilon>0$, there exists an odd function $\Phi:\mathbb{R}\to[-1,1]$ such that $\Phi(t)\to\pm 1$ as $t\to\pm\infty$ with the following properties.
  1. (1)
    For any $s\in[1,\infty)$ and all $v\in E$, $\|F_{s,v}-\Phi(B_{s,v})\|<\varepsilon$.
  2. (2)
    There exists $R_{0}>0$ such that for all $s\in[1,\infty)$, $p\geq 2$ and $v\in E$, $\textup{Prop}_{E,p}(\Phi(B_{s,v}))\leq s^{-1}R_{0}$.
  3. (3)
    For any $s\in[1,\infty)$ and all $v\in E$, the operator $\Phi(B_{s,v})^{2}-1$ is compact.
  4. (4)
    For any $s\in[1,\infty)$ and all $v,w\in E$, the operator $\Phi(B_{s,v})-\Phi(B_{s,w})$ is compact.
  5. (5)
    There exists $c>0$ such that for all $s\in[1,\infty)$, $p\geq 2$ and all $v,w\in E$, Φ(B_s,v)-Φ(B_s,w)∥≤cv-w_p.
  6. (6)
    For any $v\in E$, the function [1,)B(L^2(E,h_E))  sΦ(B_s,v) is continuous under the strong operator topology.
  7. (7)
    The family of functions [1,)B(L^2(E,h_E))  sΦ(B_s,v)^2-1 is norm equi-continuous as $v$ varies over $E$ and $s$ varies over any fixed compact subset of $1,\infty)$.
  8. (8)
    For any $r\geq 0$, the family of functions [1,)B(L^2(E,h_E))  sΦ(B_s,v)-Φ(B_s,w) is norm equi-continuous as $(v,w)$ varies over $E\times E$ with $|v-w|\leq r$, and $s$ varies over any fixed compact subset of $1,\infty)$.
  9. (9)
    There exists $R_{1}>0$ such that for all $R\geq R_{1}$, $s\in[L\cdot(\operatorname{dim}E_{n})^{2},\infty)$, and $v\in E$ we have that: ——(Φ(B_s,v)^2-1)(1-χ_v,R)——¡3ε.
  10. (10)
    For any $r\geq 0$, there exists $R_{2}>0$ such that for all $R\geq R_{2}$, $s\in[L\cdot(\operatorname{dim}E_{n})^{2},\infty)$, and $v,w\in E\times E$ with $|x-y|\leq r$ we have that: ——(Φ(B_s,v)-Φ(B_s,w))(1-χ_v,R)——¡3ε.
We remark that the choice of $R_{1}$ in (9) and the choice of $R_{2}$ in (10) do not depend on the dimension of EE.∎
3.2. Step 2. Construction of the twisted Roe algebra
The first step ensures that, without loss of generality, we may assume $X=\bigsqcup_{n=0}^{\infty}X_{n}$ is a sparse space that admits a coarse embedding into $\ell^{p}$. By [HIT2020, Lemma 12.5.4], we can assume the distance between $X_{n}$ and $X_{m}$ is infinite whenever $n\neq m$ (called separated disjoint union of $\{X_{n}\}_{n\in\mathbb{N}}$). In this case, the Hilbert space $\mathcal{H}_{X}$ has a canonical decomposition $\mathcal{H}_{X}=\bigoplus_{n\in\mathbb{N}}\mathcal{H}_{X_{n}}$. The Roe algebra $C^{*}(X)$ is therefore a subalgebra of $\prod_{n\in\mathbb{N}}C^{*}(X_{n})$
The coarse disjoint union of $\{X_{n}\}_{n\in\mathbb{N}}$ admits a coarse embedding into an $\ell^{p}$-space if and only if each $X_{n}$ admits a coarse embedding into the $\ell^{p}$-space with respect to the same controlling functions $\rho_{+},\rho_{-}:\mathbb{R}_{+}\to\mathbb{R}_{+}$. Let $f_{n}:X_{n}\to\ell^{p}$ be a coarse embedding. Then $\mathrm{span}\{f(X_{n})\}$ forms a finite-dimensional subspace of $\ell^{p}$, which must be coarsely equivalent to a finite-dimensional $\ell^{p}$ space. Without loss of generality, we may assume the image of $f_{n}$ is a finite-dimensional linear space $(E_{n},\|\cdot\|_{p})$. Set $d_{n}=\dim(E_{n})$. We shall write $E$ to be the collection of all $\{E_{n}\}$.
Define
$\mathcal{H}_{X_{n},E_{n}}=\mathcal{H}_{X_{n}}\otimes L^{2}(E_{n},\mathfrak{h}_{E_{n}})\quad\text{and}\quad\mathcal{H}_{X,E}=\bigoplus_{n\in\mathbb{N}}\mathcal{H}_{X_{n},E_{n}}.$
Note that we define the Roe algebra $C^{*}(X,\mathcal{H}_{X,E})$ with respect to $\mathcal{H}_{X,E}$. We will consider $C^{*}(X_{n},\mathcal{H}_{X_{n}})$ as represented on $\mathcal{H}_{X_{n},E_{n}}=\mathcal{H}_{X_{n}}\otimes L^{2}(E_{n},\mathfrak{h}_{E_{n}})$ via the amplified representation $T\mapsto T\otimes 1$. Thus $C^{*}(X,\mathcal{H}_{X})$ can be viewed as a subalgebra of the multiplier algebra of $C^{*}(X,\mathcal{H}_{X,E})$.
Here, $\mathcal{H}_{X,E}$ admits natural actions of $C_{0}(X)$ and $C_{0}(E)$. Hence, for an operator $T$ on this space, we can discuss its propagation in the $X$-direction and its propagation in the $E$-direction, denoted by $\textup{Prop}_{X}(T)$ and $\textup{Prop}_{E,p}(T)$, respectively.
For any $v\in E_{n}$, we set $\chi^{(n)}_{R,p,v}$ to be the characteristic function of the ball $B(v,R)$ in $(E_{n},\|\cdot\|_{p})$. Consider the function $\chi^{(n)}_{R,p}:X\mapsto B(L^{2}(E_{n},\mathfrak{h}_{E_{n}}))$ defined by
$\chi^{(n)}_{R,p}(x)=\chi^{(n)}_{R,p,f_{n}(x)}.$
Notice that each $\chi^{(n)}_{R,p,v}$ can be viewed as a multiplication operator on $L^{2}(E_{n},\mathfrak{h}_{E_{n}})$. Thus, $\chi^{(n)}_{R,p}$ can be viewed as a bounded operator on $\mathcal{H}_{X_{n},E_{n}}$.
Now, we are ready to define the twisted algebras:
Definition 3.4.
Let $\prod_{n\in\mathbb{N}}C_{b}([1,\infty),C^{*}(X_{n},\mathcal{H}_{X_{n},E_{n}}))$ be the direct product $C^{*}$-algebra of all bounded continuous functions from $1,\infty)$ to $C^{*}(X_{n},\mathcal{H}_{X_{n},E_{n}})$. Write elements of this algebra as collection $T=(T_{s,n})_{n\in\mathbb{N},s\in[1,\infty)}$. Let $\mathbb{A}[X,E]$ be the *-subalgebra of $\prod_{n\in\mathbb{N}}C_{b}([1,\infty),C^{*}(X_{n},\mathcal{H}_{X_{n},E_{n}}))$ consisting of elements satisfying the following conditions
  • (1)
    $\sup\limits_{s\in[1,\infty),n\in\mathbb{N}}\textup{Prop}_{X_{n}}(T_{s,n})<\infty$;
  • (2)
    $\lim\limits_{s\to\infty}\sup\limits_{n\in\mathbb{N}}\textup{Prop}_{E_{n},p}(T_{s,n})=0$;
  • (3)
    $\lim\limits_{R\to\infty}\sup\limits_{s\in[1,\infty),n\in\mathbb{N}}\|\chi^{(n)}_{R,p}T_{s,n}-T_{s,n}\|=\lim\limits_{R\to\infty}\sup\limits_{s\in[1,\infty),n\in\mathbb{N}}\|T_{s,n}\chi^{(n)}_{R,p}-T_{s,n}\|=0$.
Define the twisted Roe algebra $A(X,E)$ to be the closure of $\mathbb{A}[X,E]$ with respect to the norm induced from $\prod_{n\in\mathbb{N}}C_{b}([1,\infty),C^{*}(X_{n},\mathcal{H}_{X_{n},E_{n}}))$.
Next, we will use the bounded Bott-Dirac map from 3.2 to construct the index map:
$\textup{Ind}:K_{*}(C^{*}(X,\mathcal{H}_{X}))\to K_{*}(A(X,E)).$
For each $s\in[1,\infty)$, $F_{s,v,E_{n}}:L^{2}(E_{n},\mathfrak{h}_{E_{n}})\to L^{2}(E_{n},\mathfrak{h}_{E_{n}})$ is the bounded operator as in 3.2. Set $s_{n}:=L(\operatorname{dim}E_{n})^{2}$, where $L$ is the constant in 3.3.
Definition 3.5.
Let $F:[1,\infty)\to\mathcal{B}(\mathcal{H}_{X,E})$ be a function
$s\mapsto F_{s}=(F_{s,1},\cdots,F_{s,n},\cdots)\in\prod_{n\in\mathbb{N}}\mathcal{B}(\mathcal{H}_{X_{n},E_{n}})\subseteq\mathcal{B}(\mathcal{H}_{X,E}),$
where $F_{s,n}\in\mathcal{B}(\mathcal{H}_{X_{n},E_{n}})$ is defined to be
$(F_{s,n})_{xy}=\left\{\begin{aligned} &F_{s+s_{n},f_{n}(x),E_{n}},&&\text{if }x=y;\\ &0,&&\text{if }x\neq y.\end{aligned}\right$.
Here, we view each element in $\mathcal{B}(\mathcal{H}_{X_{n},E_{n}})$ as an $X_{n}$-by-$X_{n}$ matrix, where each entry of the matrix is a bounded linear operator on $L^{2}(E_{n},\mathfrak{h}_{E_{n}})$.
Analogously, if $\Phi$ has the properties in 3.3 for some $\varepsilon>0$, we write $F^{\Phi}$, $F^{\Phi}_{s}$ and $F^{\Phi}_{s,n}$ for the operators built in the same way as $F$, $F_{s}$ and $F_{s,n}$, but replacing $F_{s,v,E_{n}}$ with $\Phi(B_{s,v,E_{n}})$.
By 3.3, the function $F$ is strong *-continuous with respect to $s\in[1,\infty)$. To introduce the index map, we also need the following lemma. The proof of the lemma can be found in [WXYZ2024, Proposition 2.19] and [HIT2020, Section 12.3]. We will omit some details here.
Lemma 3.6.
  • (1)
    The function $F$ defined above is an odd, self-adjoint, contractive operator in the multiplier algebra of $A(X,E)$.
  • (2)
    The Roe algebra $C^{*}(X,H_{X})$ is a $C^{*}$-subalgebra of the multiplier algebra of $A(X,E)$.
  • (3)
    For any $T\in C^{*}(X,H_{X})$ and $s\in[1,\infty)$, the function $s\mapsto[T,F_{s}]$ and $s\mapsto T(F^{2}_{s}-1)$ is in $A(X,E)$.
  • (4)
    For any $P\in C^{*}(X,H_{X})$ a projection, the function
    $s\mapsto(PF_{s}P)^{2}-P$
    is in the corner PA(X,E)PPA(X,E)P.∎
It should be noted that in 3.5, the parameter $s$ is shifted to the right by $s_{n}$ (depending on the dimension) to ensure the validity of the preceding 3.6 applying (9), (10) in 3.3.
By [HIT2020, Corollary 2.7.4], a $K_{0}$-theory element of $C^{*}(X)$ can be represented by a single projection $[P]$. For any projection $P\in C^{*}(\mathcal{H}_{X})$, $PF_{s}P$ is an odd, self-adjoint operator on the graded Hilbert space $L^{2}[1,\infty)\otimes\mathcal{H}_{X,E}$. We can build an index class of $K_{0}(PA(X,E)P)$ by using $PF_{s}P$ and 3.6. We can decompose $PF_{s}P$ as a matrix with respect to the grading on each $\mathcal{H}_{X,E}$ and $PF_{s}P$ has the form
$PF_{s}P=\begin{pmatrix}0&u^{*}\\ u&0\end{pmatrix}$
where $P$ is even and has the form
$P=\begin{pmatrix}P_{0}&0\\ 0&P_{1}\end{pmatrix}$.
We define the index class of $PF_{s}P$ in $K_{0}(PA(X,E)P)$, denoted by $\textup{Ind}(PF_{s}P)$, to be
$\textup{Ind}(PF_{s}P)=\begin{bmatrix}(P_{0}-u^{*}u)^{2}&u^{*}(P_{1}-uu^{*})\\ u(2P_{0}-u^{*}u)(P_{0}-u^{*}u)&uu^{*}(P_{1}-uu^{*})\end{bmatrix}-\begin{bmatrix}0&0\\ 0&P_{1}\end{bmatrix}$
We define the inclusion homomorphism: $i_{P}:PA(X,E)P\to A(X,E)$. It induces a group homomorphism $(i_{P})_{*}:K_{*}(PA(X,E)P)\to K_{*}(A(X,E))$ on the level of $K$-theory.
Define the index map
$\textup{Ind}:K_{0}(C^{*}(\mathcal{H}_{X}))\to K_{0}(A(X,E))$
by the following formula:
$P]\mapsto(i_{P})_{*}(\textup{Ind}(PF_{s}P)).$
By using a suspension argument, we can define the index map on the level of $K_{1}$. Similarly, we can also define the localization index map for both $K_{0}$ and $K_{1}$ by applying the above construction pointwise in $t$.
Definition 3.7.
Define the index map as we constructed above:
$\textup{Ind}:K_{*}(C^{*}(\mathcal{H}_{X}))\to K_{*}(A(X,E)).$
The index map is a well-defined group homomorphism. For each $s\in[1,\infty)$, define $\iota^{s}:A(X,E)\to C^{*}(X,H_{X,E})$ to be the evaluation map $(T_{s})\mapsto T_{s}$. The following proposition is proved in [WXYZ2024, Proposition 2.20], also see [HIT2020, Section 12.3]. We shall provide a detailed proof in 3.12.
Proposition 3.8.
For any $s\in[1,\infty)$, the composition
$K_{*}(C^{*}(X,\mathcal{H}_{X}))\stackrel{{\scriptstyle ind}}{{\longrightarrow}}K_{*}(A(X,E))\stackrel{{\scriptstyle\iota^{s}_{*}}}{{\longrightarrow}}K_{*}(C^{*}(X,\mathcal{H}_{X,E}))$
is an isomorphism.∎
3.3. Step 3. Geometric and Ghostly ideals in the twisted setting
In this subsection, we consider restricting the above Dirac-dual-Dirac construction to the geometric ideal and the ghostly ideal.
Definition 3.9.
For any $(T_{s,n})\in A(X,E)$, we shall define the $\varepsilon$-support of $(T_{s,n})$ to be the set
$\textup{supp}_{\varepsilon}((T_{s,n}))=\bigcup_{n\in\mathbb{N}}\left\{(x,y)\in X_{n}\times X_{n}\ \Big|\ \sup_{s\in[1,\infty)}\|(T_{s,n})_{xy}\|\geq\varepsilon\right\}.$
Fix an invariant open set $U\subseteq\beta X$. The geometric ideal of $A(X,E)$ associated with $U$, denoted by $A_{\mathcal{I}(U)}(X,E)$, is defined to be the ideal generated by
$\mathbb{A}_{\mathcal{I}(U)}[X,E]=\left\{(T_{s,n})\in\mathbb{A}[X,E]\ \big|\ \overline{r(\textup{supp}_{\varepsilon}((T_{s,n})))}\subseteq U\right\}.$
The ghostly ideal of $A(X,E)$ associated with $U$, denoted by $A_{\mathcal{G}(U)}(X,E)$, is defined to be the algebra
$A_{\mathcal{G}(U)}(X,E)=\left\{(T_{s,n})\in A(X,E)\ \big|\ \overline{r(\textup{supp}_{\varepsilon}((T_{s,n})))}\subseteq U\right\}.$
Note that the distinction between the geometric ideal and the ghostly ideal in the twisted algebra here is parallel to the distinction in the original Roe algebra, which is the difference in the order of taking the $\varepsilon$-support and the completion. It is easy to show that both $A_{\mathcal{I}(U)}(X,E)$ and $A_{\mathcal{G}(U)}(X,E)$ can absorb $\mathbb{A}[X,E]$, and thus they indeed give two ideals in $A(X,E)$.
Lemma 3.10.
Let $T\in\mathcal{I}(X,U)$ (or $T\in\mathcal{G}(X,U)$, respectively). Then for any $s\in[1,\infty)$, the paths $s\mapsto[T,F_{s}]$ and $s\mapsto T(F_{s}^{2}-1)$ belong to $A_{\mathcal{I}(U)}(X,E)$ (or $A_{\mathcal{G}(U)}(X,E)$, respectively).
Proof.
We will only prove the case for the map $s\mapsto T(F_{s}^{2}-1)$ by using (7) in 3.3, as the argument for $T,F_{s}]$ is analogous by using (8) in 3.3. By 3.6, we already have these two functions belong to $A(X,E)$.
Note that every operator on $\mathcal{H}_{X,E}$ can be viewed as an $X$-by-$X$ matrix, where each entry determines an element in $B(L^{2}(E,\mathfrak{h}_{E}))$. Then, for any $s\in[1,\infty)$, we have
$(T(F_{s}^{2}-1))_{xy}=T^{(n)}_{xy}\cdot(F_{s,f(x),E_{n}}^{2}-1),$
where $x,y\in X_{n}$. Thus, we have that
$\|(T(F_{s}^{2}-1))_{xy}\|\leq\|T^{(n)}_{xy}\|\cdot\|(F_{s,f(x),E_{n}}^{2}-1)\|\leq\|T^{(n)}_{xy}\|$.
As a conclusion,
$\textup{supp}_{\varepsilon}(T(F_{s}^{2}-1))\subseteq\textup{supp}_{\varepsilon}(T).$ (1)
For the case of the geometric ideal, it suffices to show that for $T\in\mathbb{I}[X,U]$, we have $T(F_{s}^{2}-1)\in A_{\mathcal{I}(U)}[X,E]$. Then it is direct from the inclusion (1). ∎
Corollary 3.11.
For any projection $P\in\mathcal{I}(X,U)$ (or $\mathcal{G}(X,U)$, respectively), the function
$s\mapsto(PF_{s}P)^{2}-P$
is in the corner $PA_{\mathcal{I}(U)}(X,E)P$ (or $PA_{\mathcal{G}(U)}(X,E)P$, respectively).
Proof.
Notice that
$\begin{split}(PF_{s}P)^{2}-P&=P(F_{s}PF_{s}-P)P\\ &=P(F_{s}PF_{s}-F^{2}_{s}P+F^{2}_{s}P-P)P\\ &=PF_{s}[P,F_{s}]P+P(F^{2}_{s}-1)P.\end{split}$
Combining this with 3.10, we have that
$s\mapsto PF_{s}[P,F_{s}]P\quad\text{ and }\quad s\mapsto P(F^{2}_{s}-1)P$
are in $A_{\mathcal{I}(U)}(X,E)$ (or $A_{\mathcal{G}(U)}(X,E)$, respectively) if $P$ is in $\mathcal{I}(X,U)$ (or $\mathcal{G}(X,U)$, respectively). ∎
To distinguish, we denote the geometric ideal and the ghostly ideal of $C^{*}(X,\mathcal{H}_{X,E})$ as $\mathcal{I}_{E}(X,U)$ and $\mathcal{G}_{E}(X,E)$, respectively. By 3.11, we can then restrict the index map on the geometric ideal and ghostly ideal. Thus, we have the following commuting diagram
[…]
Analogous to 3.8, we also have the following proposition for ideals. We should mention that the proof of 3.12 can also be applied to 3.8.
Proposition 3.12.
For any $s\in[1,\infty)$, the composition
$K_{*}(\mathcal{I}(X,U))\stackrel{{\scriptstyle\textup{Ind}}}{{\longrightarrow}}K_{*}(A_{\mathcal{I}(U)}(X,E))\stackrel{{\scriptstyle\iota^{s}_{*}}}{{\longrightarrow}}K_{*}(\mathcal{I}_{E}(X,U))$
and
$K_{*}(\mathcal{G}(X,U))\stackrel{{\scriptstyle\textup{Ind}}}{{\longrightarrow}}K_{*}(A_{\mathcal{G}(U)}(X,E))\stackrel{{\scriptstyle\iota^{s}_{*}}}{{\longrightarrow}}K_{*}(\mathcal{G}_{E}(X,U))$
are isomorphisms.
Proof.
We shall only prove the $K_{0}$-case, and the $K_{1}$-case follows from the suspension argument. The proof for $\mathcal{I}(X,U)$ also applies for $\mathcal{G}(X,U)$, thus we shall focus on $\mathcal{I}(X,U)$.
We define the map $\kappa:(E,\|\cdot\|_{p})\to(E,\|\cdot\|_{p})$ to be
$\kappa(v)=\left\{\begin{aligned} &0&&,\|v\|_{p}\leq 1\\ &v-\frac{v}{\|v\|_{p}}&&,\text{otherwise}.\end{aligned}\right$.
To simplify the notation, for fixed $s\in[1,\infty)$, we write $F_{v,n}$ for $F_{s,v,E_{n}}$. For each $k,n\in\mathbb{N}$, we define the bounded operator $F^{(k)}_{n}:\mathcal{H}_{X_{n},E_{n}}\to\mathcal{H}_{X_{n},E_{n}}$ to be
$(F^{(k)}_{n})_{xy}=\left\{\begin{aligned} &1\otimes F_{\kappa^{k}(f_{n}(x)),n}&&,x=y\in X_{n}\\ &0&&,x\neq y.\end{aligned}\right$.
Moreover, we define $F^{(\infty)}_{n}:\mathcal{H}_{X_{n},E_{n}}\to\mathcal{H}_{X_{n},E_{n}}$ by the formula
$(F^{(\infty)}_{n})_{xy}=\left\{\begin{aligned} &1\otimes F_{0,n}&&,x=y\in X_{n}\\ &0&&,x\neq y.\end{aligned}\right$.
Notice that for each $k\in\mathbb{N}\cup\{\infty\}$, $F^{(k)}=(F^{(k)}_{n})_{n\in\mathbb{N}}$ is in the multiplier of $\mathcal{I}_{E}(X,U)$ by 3.3. Then for any projection $P\in\mathcal{I}(X,U)$, we can define an index map
$\textup{Ind}^{(k)}:K_{*}(\mathcal{I}(X,U))\to K_{*}(\mathcal{I}_{E}(X,U))$
as in Definition 3.7.
By definition, one has that $\textup{Ind}^{(0)}=\iota^{s}_{*}\circ\textup{Ind}:K_{*}(\mathcal{I}(X,U))\to K_{*}(\mathcal{I}_{E}(X,U))$. We claim that $\textup{Ind}^{(0)}=\textup{Ind}^{\infty}$. To see this, let
$\mathcal{H}^{\infty}_{X_{n},E_{n}}=\ell^{2}(X_{n})\otimes\mathcal{H}\otimes(\mathcal{L}^{2}(E_{n},\mathfrak{h}_{E_{n}}))^{\oplus\infty}\quad\text{ and }\quad H^{\infty}_{X,E}=\bigoplus_{n\in\mathbb{N}}H^{\infty}_{X_{n},E_{n}}$
The Hilbert space $H^{\infty}_{X,E}$ is also a geometric module of $X=\bigsqcup X_{n}$. Denote $C^{*}(X,H^{\infty}_{X,E})$ the corresponding Roe algebra and $\mathcal{I}_{E}^{\infty}(X,U)$ its geometric ideal. The top-left corner inclusion $\mathcal{I}(X,U)\to\mathcal{I}_{E}^{\infty}(X,U)$ defined by $T=(T_{n})\mapsto(T_{n}\oplus_{k\geq 1}0)$. It induces an isomorphism on the $K$-theory level.
For each $P]\in K_{0}(\mathcal{I}(X,U))$ [^1] By [HIT2020, Corollary 2.7.4], a $K_{0}$-theory element of $\mathcal{I}(X,U)$ can be represented by a single projection $[P]$., we write $\textup{Ind}^{(k)}([P])=[(p^{(k)}_{0},\cdots,p^{(k)}_{n},\cdots)]-[q_{0},\cdots,q_{n}\cdots]$ for each $k\in\mathbb{N}\cup\{\infty\}$, where $q_{n}$ does not depend on $k$ but only depends on $[P]$ by definition. For simplicity of notation, we write
$0\oplus_{k\geq 1}p^{(k)}_{n}=\begin{pmatrix}0&0&0&\cdots\\ 0&p^{(1)}_{n}&0&\cdots\\ 0&0&p^{(2)}_{n}&\\ \vdots&\vdots&&\ddots\end{pmatrix}\quad\text{and}\quad 0\oplus_{k\geq 1}p^{(\infty)}_{n}=\begin{pmatrix}0&0&0&\cdots\\ 0&p^{(\infty)}_{n}&0&\cdots\\ 0&0&p^{(\infty)}_{n}&\\ \vdots&\vdots&&\ddots\end{pmatrix}$
for each $n\in\mathbb{N}$. We now define the two elements in $K_{0}(\mathcal{I}_{E}^{\infty}(X,U))$
[…]
$b=[(p^{(0)}_{0}\oplus_{k\geq 1}0,\cdots,p^{(0)}_{n}\oplus_{k\geq 1}0,\cdots)]-[(p^{(\infty)}_{0}\oplus_{k\geq 1}0,\cdots,p^{(\infty)}_{n}\oplus_{k\geq 1}0,\cdots)].$
We will give some explanation of why $a\in K_{0}(\mathcal{I}_{E}^{\infty}(X,U))$ here. The propagation of $a$ is less than $5\textup{Prop}_{P}(P)$ by definition of the index map, then $a$ can be approximated by operators with finite propagation as $P$ can. To see $a$ is locally compact, for each compact set $K\subset X$, there exists $N\in\mathbb{N}$ such that $X_{n}\cap K\neq\emptyset$ for all $n>N$. Note that $f_{n}(X_{n})\subset E_{n}$ is uniformly bounded by $M\in\mathbb{N}$, i.e. $f_{n}(X_{n})\subseteq B(0,M)$ for all $n\leq N$. Then we have that $\chi_{K}p_{n}^{(k)}=\chi_{K}p^{(\infty)}_{n}$ for all $k\geq M$ and $n\leq N$ since $\kappa^{k}(f_{n}(x))=0$ in this case. It shows that $\chi_{K}a$ is a direct sum of only finite compact operators, which is also compact.
Notice that $b$ is the image of $\textup{Ind}^{(0)}([P])-\textup{Ind}^{(\infty)}([P])$ under the top-left corner inclusion. It suffices to show $b=0\in K_{*}(\mathcal{I}_{E}^{\infty}(X,U))$. Consider the path $0,1]\to B(H_{X,E})$, $t\mapsto F^{(k+t)}=(F^{(k+t)}_{n})$, where
$(F^{(k+t)}_{n})_{xy}=\left\{\begin{aligned} &1\otimes F_{(1-t)\kappa^{k}(f_{n}(x))+t\kappa^{k+1}(f_{n}(x)),n}&&,x=y\in X_{n}\\ &0&&,x\neq y.\end{aligned}\right$.
By 3.3, we have that
$\|F^{(k+t)}_{n}-F^{(k+t^{\prime})}_{n}\|\leq 3|t-t^{\prime}|$
for all $n\in\mathbb{N}$. Thus, $t\mapsto F_{n}^{(k+t)}$ is equi-continuous as $n$ varies in $\mathbb{N}$. We can built a path
$)[0,1]\to\mathcal{I}_{E}^{\infty}(X,U)\quad t\mapsto a_{t}=(\oplus_{k\in\mathbb{N}}(p^{(k+t)}_{n}))-(\oplus_{k\in\mathbb{N}}(p^{(\infty)}_{n}$))
by using the index map with respect to $F^{(k+t)}$ pointwisely, where
$\oplus_{k\in\mathbb{N}}(p^{(k+t)}_{n})=\begin{pmatrix}p^{(0+t)}&0&0&\cdots\\ 0&p^{(1+t)}_{n}&0&\cdots\\ 0&0&p^{(2+t)}_{n}&\\ \vdots&\vdots&&\ddots\end{pmatrix}$.
Notice that $a_{0}]=[a+b]\in K_{0}(\mathcal{I}_{E}^{\infty}(X,U))$ and it is direct to see $a_{1}]=[a]$ by using a rotation homotopy. Thus $[b]=0$ as claimed.
Finally, we still need to show $\textup{Ind}^{(\infty)}$ is an isomorphism. Recall that $f(x)=x(1+x^{2})^{-\frac{1}{2}}$ and $F_{s,0,E_{n}}=f(s^{-1}D+C)$. For each $n\in\mathbb{N}$, we define a homotopy $0,1]\to\mathcal{K}(L^{2}(E_{n},\mathfrak{h}_{E_{n}}))$ by
$t\mapsto P_{E_{n}}(t)=\left\{\begin{aligned} &f^{2}(t^{-1}((s+s_{n})^{-1}D_{E_{n}}+C_{E_{n}}))-1&&,t\in(0,1];\\ &p_{k}&&,t=0,\end{aligned}\right$.
where $p_{k}$ is the one-dimensional projection onto the kernel of $B$. We denote $P]=[(P_{0},\cdots,P_{n},\cdots)]$ and $\textup{Ind}^{(\infty)}([P])=[(Q_{0},\cdots,Q_{n},\cdots)]$. By using the homotopy $(P_{E_{n}}(t))_{t\in[0,1]}$, one can check that $Q_{n}$ is homotopic to $P\otimes p_{k}$. However, the homotopy is not equi-continuous as the dimension of $E_{n}$ increases. The reason comes from that we replaced the parameter $s$ by $s+s_{n}$ for each $E_{n}$. One can check that $\|F_{E_{n}}(t)-F_{E_{n}}(t^{\prime})\|\leq|f(\frac{2}{t(s+s_{n})})-f(\frac{2}{t^{\prime}(s+s_{n})})|\leq(s+s_{n})|t-t^{\prime}|$ the Lipschitz constant is depended on $n$.
However, we can use a ”stacking argument” as in [HIT2020, Proposition 12.6.3] to get around this problem. Define $0,1]\mapsto\mathcal{I}_{E}(X,U)$ by $t\mapsto P(t)=(P_{0}(t),\cdots,P_{n}(t),\cdots)$ where $P_{n}(t)$ is the homotopy between $Q$ and $P\otimes p_{k}$ for each $n\in\mathbb{N}$. Let $N_{n}=[s+2s_{n}]+1$. For each $t\in[0,1]$, we define $P^{N}(t)=(P^{N}_{n}(t))\in\mathcal{I}_{E}(X,U)$
$P^{N}_{n}(t)=\begin{pmatrix}P_{n}(0)&0&0&\cdots\\ 0&P_{n}(\frac{1}{N_{n}})&0&\cdots\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&P_{n}(1)\end{pmatrix}-\begin{pmatrix}0&0&0&\cdots\\ 0&P_{n}(\frac{t}{N_{n}})&0&\cdots\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&P_{n}(\frac{N_{n}-1}{N_{n}}+\frac{t}{N_{n}})\end{pmatrix}$.
Notice that $\|F_{E_{n}}(\frac{t}{N_{n}})-F_{E_{n}}(\frac{t^{\prime}}{N_{n}})\|\leq\frac{s+s_{n}}{N_{n}}|t-t^{\prime}|\leq|t-t^{\prime}|$. Thus $P^{N}(t)$ is norm-continuous. Note that $P^{N}(0)]=[P(0)]$ and we also have $P^{N}(1)]=[P(1)]$ by a rotation homotopy. Thus, we have
$\textup{Ind}^{(\infty)}([P])=[P\otimes p_{k}]$
which is an isomorphism. ∎
3.4. Step 4. Twisted geometric ideal vs. twisted ghostly ideal
In the final part of the proof, we show that in the twisted algebra, ghostly elements automatically lie in the geometric ideal; that is, the operations of taking the completion and taking the $\varepsilon$-support commute.
Lemma 3.13.
For any $(T_{s,n})\in A_{\mathcal{G}(U)}(X,E)$ and $\varepsilon>0$, there exists $(T^{\prime}_{s,n})\in\mathbb{A}_{\mathcal{I}(U)}[X,E]$ and $S>0$ such that
$\sup_{s\in[S,\infty),n\in\mathbb{N}}\|T_{s,n}-T^{\prime}_{s,n}\|\leq\varepsilon.$
Proof.
By 3.4, we have that
$\lim\limits_{R\to\infty}\sup\limits_{s\in[1,\infty),n\in\mathbb{N}}\|\chi^{(n)}_{R,p}T_{s,n}-T_{s,n}\|=\lim\limits_{R\to\infty}\sup\limits_{s\in[1,\infty),n\in\mathbb{N}}\|T_{s,n}\chi^{(n)}_{R,p}-T_{s,n}\|=0.$
For any $\varepsilon>0$, there exists $R>0$ such that
$\sup\limits_{s\in[1,\infty),n\in\mathbb{N}}\|\chi^{(n)}_{R,p}T_{s,n}-T_{s,n}\|\leq\frac{\varepsilon}{2}\quad\text{and}\quad\sup\limits_{s\in[1,\infty),n\in\mathbb{N}}\|T_{s,n}\chi^{(n)}_{R,p}-T_{s,n}\|\leq\frac{\varepsilon}{2}.$
Without loss of generality, we shall replace $T_{s,n}$ by $\chi^{(n)}_{R,p}T_{s,n}\chi^{(n)}_{R,p}$ as the norm difference between them is at most $\varepsilon$. In this case, for any $x,y\in X_{n}$, the entry $(T_{s,n})_{xy}$ defines an element in $\mathcal{K}(L^{2}(E_{n},\mathfrak{h}_{E_{n}})\otimes\mathcal{H})$ and
$\chi^{(n)}_{R,f(x),p}(T_{s,n})_{xy}\chi^{(n)}_{R,f(y),p}=(T_{s,n})_{xy}.$ (3)
Choose a sufficiently large $S>0$ such that $\sup_{s\geq S,n}\textup{Prop}_{E,p}(T_{s,n})<R$. Equivalently, we have
$\sup_{x,y\in X,s\geq S,n\in\mathbb{N}}\textup{Prop}_{E,p}((T_{s,n})_{xy})<R.$
Since each entry is an element in $\mathcal{K}(L^{2}(E_{n},\mathfrak{h}_{E_{n}})\otimes\mathcal{H})$, we can therefore discuss their propagation with respect to $(E,\|\cdot\|_{p})$. Since $f$ is a coarse embedding, there exists $M>0$ such that $\|f(x)-f(y)\|_{p}>3R$ whenever $d(x,y)>M$. Thus, whenever $d(x,y)>M$ and $s>S$, we have that
$(T_{s,n})_{xy}=\chi^{(n)}_{R,f(x),p}(T_{s,n})_{xy}\chi^{(n)}_{R,f(y),p}=\chi^{(n)}_{R,f(x),p}\cdot\chi^{(n)}_{2R,f(y),p}(T_{s,n})_{xy}\chi^{(n)}_{R,f(y),p}$.
Since $\|f(x)-f(y)\|_{p}>3R$, we have that $B(f(x),R)\cap B(f(y),R)=\emptyset$. Thus, $\chi^{(n)}_{R,f(x),p}\cdot\chi^{(n)}_{2R,f(y),p}=0$. This means that $(T_{s,n})_{xy}=0$ whenever $d(x,y)>M$ and $s>S$, i.e., $T_{s,n}$ has finite $X$-propagation for large $s$.
Define
$T^{\prime}_{s,n}=\left\{\begin{aligned} &T_{s,n},&&\text{when }s\geq S;\\ &T_{S,n},&&\text{when }s\leq S.\end{aligned}\right$.
That is, $T^{\prime}_{s,n}$ coincides with $T_{s,n}$ for $s>S$, and for $s\leq S$, we extend it as a constant function to define it on the entire interval $1,\infty)$. Note that $T^{\prime}_{s,n}\in\mathbb{A}[X,E]$ since it has finite $X$-propagation for all $s$, and by definition, we also have $T^{\prime}_{s,n}\in A_{\mathcal{G}(U)}(X,E)$. Directly from the definition, it is easy to verify that $\mathbb{A}[X,E]\cap A_{\mathcal{G}(U)}(X,E)=\mathbb{A}_{\mathcal{I}(U)}[X,E]$. Therefore, $T^{\prime}_{s,n}$ satisfies the requirements of the lemma. ∎
As a conclusion, we have the following result.
Corollary 3.14.
The canonical inclusion $i:A_{\mathcal{I}(U)}(X,E)\to A_{\mathcal{G}(U)}(X,E)$ induces an isomorphism on $K$-theory, i.e.,
$i_{*}:K_{*}(A_{\mathcal{I}(U)}(X,E))\to K_{*}(A_{\mathcal{G}(U)}(X,E)).$
Proof.
Since the K-theory of $A_{\mathcal{I}(U)}(X,E)$ and $A_{\mathcal{G}(U)}(X,E)$ only depends on their asymptotic behavior as $s\to\infty$, [HIT2020, Lemma 6.4.11], and 3.13 shows that these two algebras have the same asymptotic behavior, the conclusion thus follows naturally. We leave the details to the reader. ∎
We can now prove 2.6.
Proof of 2.6.
The theorem follows directly from the diagram (2), 3.12, 3.13 and a diagram chasing argument. ∎
Remark 3.15.
For $p\in[1,\infty)$. Let $L^{p}=L^{p}[0,1]$ denote the $L^{p}$-space on $([0,1],m)$ equipped with the canonical Lebesgue measure $m$. By [Banachspace, Proposition III.A.1], any separable $L^{p}$-space $L^{p}(\Omega,\mu)$ is isometric to the $p$-direct sum of $L^{p}$ and $\ell^{p}$. Thus, $L^{p}[0,1]$ serves as one of the most representative $L^{p}$-spaces.
We consider the standard filtration of finite-dimensional subspaces given by dyadic step functions. For each $n\in\mathbb{N}$, let $\mathcal{D}_{n}=\{[k2^{-n},(k+1)2^{-n}):0\leq k<2^{n}\}$ be the partition of $[0,1]$ into dyadic intervals. Define $W_{n}$ to be the subspace spanned by the characteristic functions of intervals in $\mathcal{D}_{n}$. Then $W_{n}$ is linearly isometric to $\ell^{p}_{2^{n}}$, and the union $W=\bigcup_{n}W_{n}$ is dense in $L^{p}$.
Suppose that $X=\bigsqcup_{n}X_{n}$ is a sparse space that admits a uniformly coarse embedding $\{f_{n}:X_{n}\to L^{p}\}$. Since $W$ is dense in $L^{p}$, without loss of generality, we may assume that $X$ coarsely embeds into $W=\bigcup_{n\in\mathbb{N}}W_{n}$. Since $X_{n}$ is a finite set, its image $f_{n}(X_{n})$ is a finite subset of $W$, there exists a sufficiently large integer $k_{n}$ such that $f_{n}(X_{n})\subseteq W_{k_{n}}$. Identifying $W_{k_{n}}$ with the subspace of $\ell^{p}$ supported on the first $2^{k_{n}}$ coordinates, we see that $f_{n}$ effectively embeds $X_{n}$ into $\ell^{p}$. Thus, the existence of a coarse embedding into $L^{p}$ implies the existence of a coarse embedding into $\ell^{p}$. By [Banachspace, Proposition III.A.1], any separable $L^{p}(\Omega,\mu)$ is isometric to the direct sum of $L^{p}[0,1]$ and $\ell^{p}$. Thus, our main result holds if we replace $\ell^{p}$ by any general separable $L^{p}$-spaces.
3.15 leads us to the following theorem:
Theorem 3.16.
The conclusion of Theorem 2.6 holds true if the assumption of coarse embeddability into $\ell^{p}$ is replaced by coarse embeddability into a $L^{p}$-space ($1\leq p<\infty$).∎
Remark 3.17.
We should mention that 3.16 can also be proved by the finite representability of $L^{p}$ in $\ell^{p}$. see [Banachspace, Section III.E.15]. Recall that a Banach space $X$ is finitely representable in $Y$ if any finite-dimensional subspace of $X$ is $(1+\varepsilon)$-isomorphic to a subspace of $Y$ for any $\varepsilon>0$.
As a consequence of [Banachspace, Section III.E.15], the family of all finite-dimensional subspaces of $L^{p}$ admits a uniformly coarse embedding into $\ell^{p}$. Thus, a sparse space $X$ admits a coarse embedding into $L^{p}$ implies that $X$ admits a coarse embedding into $\ell^{p}$.
4. Applications to coarse Baum-Connes conjectures and $ONL_{\mathcal{P}_{Fin}}$
In this section, we present some corollaries of our main theorem and some remarks.
4.1. Applications
In [GWZ2025], a relative version of the coarse Baum-Connes conjecture was introduced. In fact, our main result has the following corollary:
Corollary 4.1.
Let $X$ be a metric space with bounded geometry that admits a coarse embedding into an $\ell^{p}$-space ($p\geq 1$). For any subspace $Y$, the relative coarse Baum-Connes conjecture holds for $(X,Y)$. In particular, the boundary coarse Baum-Connes conjecture holds for $X$.
For the convenience of the reader, we briefly recall the definition of the relative coarse Baum-Connes conjecture. Let $X$ be a metric space with bounded geometry and let $Y\subseteq X$ be a subspace. Let $C^{*}(X)$ be the Roe algebra of $X$ and $\mathcal{G}(X,U_{Y})$ be the ghostly ideal generated by $Y$. We define the relative Roe algebra of the pair $(X,Y)$ as the quotient algebra
$C^{*}_{Y,\infty}(X)=C^{*}(X)/\mathcal{G}(X,U_{Y}).$
For $d\geq 0$, we define $P_{d}(X)$ as the Rips complex of $X$ at scale $d$. Its metric is defined as the hemispherical metric given in [HIT2020, Definition 7.2.8]. Let $C^{*}_{L}(P_{d}(X))$ be the localization algebra associated with $P_{d}(X)$. Define $C^{*}_{L,Y}(P_{d}(X))$ as the subalgebra of $C^{*}_{L}(P_{d}(X))$ consisting of all functions taking values in $\mathcal{G}(P_{d}(X),U_{Y})$. This is naturally an ideal. We consider the quotient
$C^{*}_{L,Y,\infty}(P_{d}(X)):=C^{*}_{L}(P_{d}(X))/C^{*}_{L,Y}(P_{d}(X)).$
Then the evaluation map at zero,
$\text{ev}:C^{*}_{L,Y,\infty}(P_{d}(X))\to C^{*}_{Y,\infty}(P_{d}(X)),\quad f\mapsto f(0),$
defines an index map
$\text{ev}_{*}:K_{*}(C^{*}_{L,Y,\infty}(P_{d}(X)))\to K_{*}(C^{*}_{Y,\infty}(P_{d}(X))).$
By 2.8, the right-hand side is a coarse equivalence invariant, and thus is isomorphic to $K_{*}(C^{*}_{Y,\infty}(X))$. The relative assembly map for $(X,Y)$ is defined as the inductive limit of these index maps as $d\to\infty$:
$\mu:\lim_{d\to\infty}K_{*}(C^{*}_{L,Y,\infty}(P_{d}(X)))\to K_{*}(C^{*}_{Y,\infty}(X)).$
If the subspace $Y$ is chosen to be a bounded set, then the relative coarse Baum-Connes conjecture is called the boundary coarse Baum-Connes conjecture, see .
Proof of 4.1.
Consider the following exact sequence:
$K_{i}(A_{Y,L,d})$$K_{i}(A_{L,d})$$K_{i}(A_{L,Y,\infty,d})$$K_{i+1}(A_{Y,L,d})$$K_{i+1}(A_{L,d})$$K_{i}(A_{Y})$$K_{i}(A)$$K_{i}(A_{Y,\infty})$$K_{i+1}(A_{Y})$$K_{i+1}(A)$$\scriptstyle{(1)}$$\scriptstyle{(2)}$$\scriptstyle{(3)}$$\scriptstyle{(4)}$$\scriptstyle{(5)}$
where $A_{Y,L,d}=C^{*}_{L,Y}(P_{d}(X))$, $A_{L,d}=C^{*}_{L}(P_{d}(X))$, $A_{L,Y,\infty,d}=C^{*}_{L,Y,\infty}(P_{d}(X))$, $A_{Y}=\mathcal{G}(X,U_{Y})$, $A=C^{*}(X)$, and $A_{Y,\infty}=C^{*}_{Y,\infty}(X)$. Since $X$ admits a coarse embedding into $\ell^{p}$, by [WXYZ2024], we have that the maps (2) and (5) are isomorphism for both $i=0,1$. By 2.6, $K_{*}(\mathcal{G}(X,U_{Y}))\cong K_{*}(\mathcal{I}(X,U_{Y}))\cong K_{*}(C^{*}(Y))$. Thus, the maps (1) and (4) are identified with the coarse Baum-Connes assembly map for $Y$, which are also isomorphisms since $Y$ also admits a coarse embedding into $\ell^{p}$. Then the corollary follows from the Five Lemma. ∎
The following definition is introduced by M. Braga, I. Farah, and A. Vignati in [BFV2024].
Definition 4.2 ([BFV2024]).
  • (1)
    For any $C^{*}$-algebra $A$, we shall denote $\mathcal{P}_{Fin}(A)$ the set of all finite rank projections in $A$.
  • (2)
    Let $\mathcal{I}$ be any index set. The uniform direct product of $C^{*}_{u}(X)$, denoted by $\prod^{u}_{\mathcal{I}}C^{*}_{u}(X)$ is the subalgebra of $\prod_{\mathcal{I}}C^{*}_{u}(X)$ generated by all families $(T_{i})_{i\in\mathcal{I}}$ with $\sup_{i\in\mathcal{I}}\textup{Prop}(T_{i})<\infty$. One can similarly define the uniform direct product of $C^{*}(X)$.
  • (3)
    The space $X$ is said to have operator norm localization for equi-approximable finite-rank projections, (abbrev. $ONL_{\mathcal{P}_{Fin}}$), if for any $\varepsilon>0$, index set $\mathcal{I}$, and element $(T_{i})\in\mathcal{P}_{Fin}(\prod^{u}_{\mathcal{I}}C^{*}_{u}(X))$, there exists $S>0$ such that for any $i\in\mathcal{I}$, there exists $\xi\in\ell^{2}(X)$ satisfying that
    $\textup{diam}(\textup{supp}(\xi))\leq S\quad\text{and}\quad\|T_{i}\xi\|\geq(1-\varepsilon)\|T_{i}\|$
  • (4)
    If we replace $\mathcal{P}_{Fin}(\prod^{u}_{\mathcal{I}}C^{*}_{u}(X))$ by $\prod^{u}_{\mathcal{I}}C^{*}_{u}(X)$, then we shall say $X$ has operator norm localization, (abbrev. ONL, see [CTWY2008]).
We also have the following corollary.
Corollary 4.3.
If $X$ admits a coarse embedding into $\ell^{p}$, then $X$ has $ONL_{\mathcal{P}_{Fin}}$.
To prove 4.3, we need the following lemma.
Lemma 4.4.
Let $X$ be a sparse metric space. All ghost projections in $C^{*}(X)$ are compact if and only if the canonical inclusion $i:\mathcal{K}\to\mathcal{G}$ induces an isomorphism $i_{*}:K_{0}(\mathcal{K})\to K_{0}(\mathcal{G})$, where $\mathcal{G}=\mathcal{G}(X,X)$ is the ghostly ideal associated with the invariant set $X\subseteq\beta X$.
Proof.
The $(\Rightarrow)$ part. For any projection $P\in M_{n}(\mathcal{G})$, since the Roe algebra is quasi-stable, there exists a projection $P^{\prime}\in\mathcal{G}$ such that $P]=[P^{\prime}]\in K_{0}(\mathcal{G})$. By assumption, $P^{\prime}$ is compact, thus $P^{\prime}]\in K_{0}(\mathcal{K})$ and $i_{*}([P^{\prime}])=[P]$. Thus $i_{*}$ is surjective. Since $X$ is sparse, the map $i_{*}$ is injective by [OOY2009, Proposition 2.10 & Remark 2.12]. This shows that $i_{*}$ is an isomorphism.
The $(\Leftarrow)$ part. Assume for a contradicition that $P\in C^{*}(X)$ is a non-compact ghost projection. Say $X=\bigsqcup_{n\in\mathbb{N}}X_{n}$ and $N_{n}=\#X_{n}$. Notice that $\prod^{u}_{n\in\mathbb{N}}C^{*}(X_{n})$ is a subalgebra of $C^{*}(X)$ and $\mathcal{K}(\mathcal{H}_{X})\cap\prod^{u}_{n\in\mathbb{N}}C^{*}(X_{n})=\bigoplus_{n\in\mathbb{N}}M_{N_{n}}(\mathcal{K})$. Moreover, the quotient algebras are canonically isomorphic to each other, i.e.,
$\frac{C^{*}(X)}{\mathcal{K}(\mathcal{H}_{X})}\cong\frac{\prod^{u}_{n\in\mathbb{N}}C^{*}(X_{n})}{\bigoplus_{n\in\mathbb{N}}M_{N_{n}}(\mathcal{K})}.$
Since each $X_{n}$ is finite, the Roe algebra $C^{*}(X_{n})$ is the algebra of compact operators. Thus, we can define $\tau:C^{*}(X)\to\prod_{n\in\mathbb{N}}\mathcal{K}/\bigoplus_{n\in\mathbb{N}}\mathcal{K}$ via the composition
$C^{*}(X)\xrightarrow{\pi}\frac{C^{*}(X)}{\mathcal{K}(\mathcal{H}_{X})}\cong\frac{\prod^{u}_{n\in\mathbb{N}}C^{*}(X_{n})}{\bigoplus_{n\in\mathbb{N}}M_{N_{n}}(\mathcal{K})}\hookrightarrow\frac{\prod_{n\in\mathbb{N}}\mathcal{K}}{\bigoplus_{n\in\mathbb{N}}\mathcal{K}}.$
This leads to a “trace map”
$\tau_{*}:K_{0}(C^{*}(X))\to K_{0}\left(\frac{\prod_{n\in\mathbb{N}}\mathcal{K}}{\bigoplus_{n\in\mathbb{N}}\mathcal{K}}\right)\cong\frac{\prod_{n\in\mathbb{N}}\mathbb{Z}}{\bigoplus_{n\in\mathbb{N}}\mathbb{Z}}.$
For any $T\in C^{*}(X)$, it is direct to see that $T$ is compact if and only if $\tau(T)=0$. For the non-compact ghost projection $P$, the $\tau(P)$ defines a non-zero projection on $\prod_{n\in\mathbb{N}}\mathcal{K}/\bigoplus_{n\in\mathbb{N}}\mathcal{K}$. Write $\tau(P)=[p_{1},\cdots,p_{n},\cdots]$ with each $p_{i}$ is a projection in $\mathcal{K}$. Since $\tau(P)$ is non-zero, we have that $\limsup_{n\to\infty}\|p_{n}\|=1$. Thus, $\tau_{*}([P])$ is non-zero.
Consider the following commuting diagram
$K_{0}(\mathcal{K})$$\frac{\prod_{n\in\mathbb{N}}\mathbb{Z}}{\bigoplus_{n\in\mathbb{N}}\mathbb{Z}}$$K_{0}(\mathcal{G})$$\scriptstyle{i_{*}}$$\scriptstyle{\tau_{*}}$$\scriptstyle{\tau_{*}}$
where $\mathcal{G}=\mathcal{G}(X,X)$ is the ghostly ideal associated with the invariant set $X\subseteq\beta X$. If $X$ admits a non-compact ghost $P\in C^{*}(X)$, then $P]\in K_{0}(\mathcal{G})$ can never be in the image of $i_{*}$ since $\tau_{*}$ maps $K_{0}(\mathcal{K})$ to 0 while $\tau_{*}([P])\neq 0$. This leads to a contradiction to the fact that $i_{*}$ is an isomorphism. ∎
Proof of 4.3.
Assume for a contradiction that $X$ does not have $ONL_{\mathcal{P}_{Fin}}$. By [BFV2024, Theorem 1.3], there exists a sparse subspace $X^{\prime}\subseteq X$ such that there exists a non-compact ghost projection in $C^{*}_{u}(X^{\prime})$. Without loss of generality, we may assume that $X$ itself is sparse and and $P\in C^{*}_{u}(X)$ is a non-compact ghost projection. Take $p\in\mathcal{K}(\mathcal{H})$ to be a rank one-projection. Then the tensor product $P\otimes p$ defines a non-compact ghost projection in $C^{*}(X)$. By 4.4, the map $i_{*}$ is not an isomorphism on $K_{0}$. This leads to a contradiction to 2.6. ∎
We should mention that 4.3 provides a large class of examples satisfying the property $ONL_{\mathcal{P}{Fin}}$, including all metric spaces that admit a coarse embedding into Hilbert space.
4.2. A remark on maximal coarse Baum-Connes conjecture
In this section, we discuss how the method used in [HIT2020] to prove the coarse Baum-Connes conjecture also applies to the maximal coarse Baum-Connes conjecture. This requires the following proposition.
Proposition 4.5.
Let $\pi:\mathbb{C}[X]\to\mathcal{B}(\mathcal{H})$ be a non-degenerated *-representation. Then there exists a representation $\rho:\ell^{\infty}(X,\mathcal{B})\to\mathcal{B}(\mathcal{H})$ on the same Hilbert space $\mathcal{H}$ such that
$\rho(\ell^{\infty}(X,\mathcal{B}))\subseteq M(C^{*}_{\pi}(X)),$
where $C^{*}_{\pi}(X)$ is the $C^{*}$-completion of $\pi(\mathbb{C}[X])$ in $\mathcal{B}(\mathcal{H})$ and $M(C^{*}_{\pi}(X))$ is the multiplier algebra of $C^{*}_{\pi}(X)$.
Proof.
Notice that $\ell^{\infty}(X,\mathcal{K})\subseteq\mathbb{C}[X]$ is a subalgebra. Moreover $\ell^{\infty}(X,\mathcal{B})$ is identified with the multiplier algebra of $\ell^{\infty}(X,\mathcal{K})$. Any *-representation of a $C^{*}$-algebra $A$ can be lift to a representation of $M(A)$, see [HIT2020, Proposition 1.7.3]. We shall denote $\rho:\ell^{\infty}(X,\mathcal{B})\to\mathcal{B}(H)$ the lifted representation of $\pi|_{\ell^{\infty}(X,\mathcal{K})}$.
By [GWY2008, Lemma 3.4], any element $T\in\mathbb{C}[X]$ can be written in the form
$T=\sum_{i=1}^{N}f_{i}V_{i}$
with $f_{i}\in\ell^{\infty}(X,\mathcal{K})$ and $V_{i}$ satisfies that $V^{*}_{i}V_{i}\in\ell^{\infty}(X,\mathcal{K})$. Thus, for any $g\in\ell^{\infty}(X,\mathcal{B})$, we have that
$\rho(g)\pi(T)=\rho(g)\sum\pi(f_{i})\pi(V_{i})=\sum\pi(gf_{i})\pi(V_{i})\in\pi(\mathbb{C}[X]).$
This means that $\rho(\ell^{\infty}(X,\mathcal{B}))\subseteq M(C^{*}_{\pi}(X))$. ∎
Lemma 4.6.
Let $X=\bigsqcup_{n\in\mathbb{N}}X_{n}$ be the separated disjoint union of finite spaces $\{X_{n}\}_{n\in\mathbb{N}}$. For any representation $\pi:\ell^{\infty}(X)\to\mathcal{B}(H)$, we have that $\sum_{n\in\mathbb{N}}\pi(\chi_{X_{n}})$ SOT-converges to $I$.
Proof.
For any $\xi\in H$, by Riesz’ representation theorem, the function $\mu:\mathcal{P}(X)\to[0,\|\xi\|]$
$\mu(E)=\langle\chi_{E}\xi,\xi\rangle$
defines a measure on $X$. Thus,
$\mu(X)=\mu(\bigsqcup_{n\in\mathbb{N}}X_{n})=\sum_{n\in\mathbb{N}}\mu(X_{n}).$
As a result, $\|\sum_{n\geq N}\pi(\chi_{X_{n}})\xi\|^{2}=\sum_{n\geq N}\mu(X_{n})$ tends to 0 as $N$ tends to infinity. ∎
By the above proposition, we can apply the methods from [HIT2020, Chapter 12] and [WXYZ2024] to the maximal coarse Baum-Connes conjecture. First, consider a maximal representation $\pi$ of the Roe algebra $\mathbb{C}[X,\mathcal{H}_{X}]$. By 4.5, since $\mathbb{C}[X,\mathcal{H}_{X}]$ and $\mathbb{C}[X,\mathcal{H}_{X,E}]$ are unitarily equivalent, this induces a maximal representation $\pi_{E}$ of $\mathbb{C}[X,\mathcal{H}_{X,E}]$. One can identify both $\ell^{\infty}(X)$ and $C_{0}(E)$ as an subalgebra of $\ell^{\infty}(X,\mathcal{B}(L^{2}(E,\mathfrak{h}_{E})\otimes\mathcal{H}))$. Therefore, even in the maximal Roe algebra, we can still define the propagation with respect to $X$ and $E$. By 4.6, we can write an element $T\in C^{*}_{\max}(X,\mathcal{H}_{X,E})$ as a sequence of $T_{n}$ with each $T_{n}=\chi_{X_{n}}T\chi_{X_{n}}$. Moreover, both $\chi_{R,p,v}$ in 3.4 and $F_{s,n}$ in 3.5 are elements in $\ell^{\infty}(X,\mathcal{B}(L^{2}(E,\mathfrak{h}_{E})\otimes\mathcal{H}))$, the condition (2) in 3.4 remains well-defined for the maximal Roe algebra.
Recalling the proof of 3.6 in [HIT2020, Chapter 12], one finds that the proofs of these statements are, in fact, carried out before completion, namely, within the algebras $\mathbb{C}[X,\mathcal{H}_{X}]$ and $\mathbb{A}[X,E]$. For example, $F$ is a multiplier of $\mathbb{A}[X,E]$, $\mathbb{C}[X,\mathcal{H}_{X}]$ is a subalgebra of the multiplier algebra of $\mathbb{A}[X,E]$, and so on. These statements therefore extend to the maximal completions by 4.5. Consequently, we can still define the index map from the maximal Roe algebra to the maximal version of the twisted algebra. Moreover, one can verify that the proofs in [HIT2020, Section 12.3 & 12.4] also hold within the dense subalgebra $\mathbb{C}[X,\mathcal{H}_{X,E}]$, and thus they naturally extend to the maximal completions. Hence, combining these observations, we obtain:
Theorem 4.7.
Let $X$ be a metric space with bounded geometry. If $X$ admits a coarse embedding into an $\ell^{p}$-space, then the maximal coarse Baum-Connes conjecture holds for XX.∎
As an application, we have the following result.
Corollary 4.8.
Let $X$ be a metric space with bounded geometry that admits a coarse embedding into an $\ell^{p}$-space ($p\geq 1$).
  • (1)
    The canonical quotient map $\pi:C^{*}_{\max}(X)\to C^{*}(X)$ induces an isomorphism on $K$-theory, i.e.,
    $\pi_{*}:K_{*}(C^{*}_{\max}(X))\xrightarrow{\cong}K_{*}(C^{*}(X)).$
  • (2)
    For any subspace $Y$, the canonical quotient map $\pi:C^{*}_{\max,Y,\infty}(X)\to C^{*}_{Y,\infty}(X)$ induces an isomorphism on the level of $K$-theory, i.e.,
    $\pi_{*}:K_{*}(C^{*}_{\max,Y,\infty}(X))\xrightarrow{\cong}K_{*}(C^{*}_{Y,\infty}(X)).$
  • (3)
    For any subspace $Y$, the maximal relative coarse Baum-Connes conjecture holds for $(X,Y)$.
Proof.
(1) Consider the following commuting diagram
$\lim\limits_{d\to\infty}K_{*}(C^{*}_{L,\max}(P_{d}(X)))$$K_{*}(C^{*}_{\max}(X))$$\lim\limits_{d\to\infty}K_{*}(C^{*}_{L}(P_{d}(X)))$$K_{*}(C^{*}(X))$$\scriptstyle{\mu_{\max}}$$\scriptstyle{(\pi_{L})*}$$\scriptstyle{\pi_{*}}$$\scriptstyle{\mu}$
The left vertical map is induced by the canonical quotient map on localization algebras
$\pi_{L}:C^{*}_{L,\max}(P_{d}(X))\to C^{*}_{L}(P_{d}(X))$
by taking quotient pointwisely. Since both maximal and reduce coarse Baum-Connes conjecture holds for $X$ (by 4.7 and [WXYZ2024, Theorem 1.1]), the maps $\mu_{\max}$ and $\mu$ are isomorphisms. Thus, it suffices to show that $\pi_{L}$ induces an isomorphism on $K$-theory, this holds from a cutting and pasting argument in [Yu1997, Theorem 3.2].
(2) Since the left-hand side of the coarse Baum-Connes conjecture coincides for the maximal and reduced norms, 4.7 implies that the map $K_{*}(C^{*}_{\max}(X))\to K_{*}(C^{*}(X))$ is an isomorphism. We then have the following commutative diagram,
[…]
where $C^{*}_{\max}(X,Y)$ denotes the geometric ideal in the maximal Roe algebra. Its $K$-theory is isomorphic to $K_{*}(C^{*}_{\max}(Y))$; see [GWZ2025] for the associated short exact sequence. For the proof of this isomorphism, one may refer to [HRY1993, Lemma 1] for the reduced case. The argument for the maximal case follows verbatim.
The left vertical map is given by the compostion
$C^{*}_{\max}(X,Y)\xrightarrow[\text{canonical quotient}]{\pi}\mathcal{I}(X,U_{Y})\xrightarrow[\text{canonical inclusion}]{i}\mathcal{G}(X,U_{Y}).$
By item (1), $K(C^{*}_{\max}(Y))$ is isomorphic to $K_{*}(\mathcal{I}(X,U_{Y}))$ since $Y$ coarsely embeds into $\ell^{p}$-space, which is further isomorphic to $K_{*}(\mathcal{G}(X,U_{Y}))$ by 2.6. Then it follows immediately from the Five Lemma.
(3) It follows directly from (2) and 4.1. ∎